### v1

# Zeta function and entropy for non-archimedean subhyperbolic dynamics

Abstract.
Let $K$ be a complete non-archimedean field equipped with a discrete valuation. We establish the rationality of the Artin-Mazur zeta function on the Julia set for any subhyperbolic rational map defined over $K$ with a compact Julia set. Furthermore, we conclude that the topological entropy on the Julia set of such a map is given by the logarithm of a weak Perron number. Conversely, we construct a (sub)hyperbolic rational map defined over $K$ with compact Julia set whose topological entropy on the Julia set equals the logarithm of a given weak Perron number. This extends Thurston’s work on the entropy for postcritically finite interval self-maps to the non-archimedean setting.
1. Introduction
The Artin-Mazur zeta function and topological entropy are deeply intertwined in the study of dynamical systems. This paper aims to investigate these two fundamental objects within the framework of non-archimedean dynamics. Specifically, we focus on systems generated by the iterative action of certain rational maps defined over non-archimedean fields.
The Artin-Mazur zeta function, introduced by Artin and Mazur [1], encodes information about the cardinality of periodic points. Its radius of convergence provides a meaningful reflection of the topological entropy for the system. In the setting of differential dynamics, the (ir)rationality of the Artin-Mazur zeta function has been extensively studied. Smale conjectured, and Manning later proved, that the Artin-Mazur zeta function is rational when the dynamics satisfies the hyperbolicity condition known as Axiom A [22]. In the context of complex dynamics on the Riemann sphere, Hinkkanen [16] established the rationality of the Artin-Mazur zeta function for rational maps. In real dynamics, Milnor and Thurston [23] studied the Artin-Mazur zeta function for interval maps, see also [25]. Meanwhile, the topological entropy has been well understood for complex rational maps [15] and thoroughly examined in the context of real dynamics [23]. Notably, Thurston [30] provided a complete description of the entropy for postcritically finite interval maps.
Non-archimedean dynamics arises from various contexts, including number theory, degenerations of complex or real dynamical systems, physics, information theory, and even cognitive science and psychology. In the setting of non-archimedean dynamics, when the ground field is algebraically closed, the study of the Artin-Mazur zeta function and the topological entropy has already appeared in [2, 8, 13, 14, 20, 29]. In this paper, we focus on univariate rational maps of degree at least $2$ defined over a non-archimedean field. Typically, in our case the ground non-archimedean field is not algebraically closed.
Throughout this paper, we fix a field $K$ which is complete with respect to a non-archimedean absolute value $|\cdot|_{K}$. Let $\phi(z)\in K(z)$ be a rational function of degree $d\geq 2$. Then $\phi$ induces a self-map on the projective line $\mathbb{P}^{1}(K)$. The $K$-Fatou set of $\phi$, denoted by $\mathcal{F}(\phi,K)$, is the maximal domain of equicontinuity of $\{\phi^{\circ n}\}_{n\geq 1}$ on $\mathbb{P}^{1}(K)$. The $K$-Julia set is then defined by $\mathcal{J}(\phi,K)=\mathbb{P}^{1}(K)\backslash\mathcal{F}(\phi,K)$. Here $\phi^{\circ n}$ denotes the $n$-times composition of $\phi$. Unlike in complex dynamics, the Julia set $\mathcal{J}(\phi,K)$ could be empty; for instance, if $\phi$ has good reduction, then $\phi$ has no Julia set for any extension of $K$, see [28, Theorem 2.17]. Moreover, by definition, $\mathcal{J}(\phi,K)$ is forward invariant under $\phi$, that is, $\phi(\mathcal{J}(\phi,K))\subseteq\mathcal{J}(\phi,K)$, and hence $\phi$ induces a dynamical system on $\mathcal{J}(\phi,K)$.
Note that any periodic point in $\mathcal{J}(\phi,K)$ must necessarily be a repelling periodic point for $\phi$. In our analysis, we focus on the Artin-Mazur zeta function $\zeta_{\phi}(K,t)$ associated with the dynamical system $(\mathcal{J}(\phi,K),\phi)$. That is,
$\zeta_{\phi}(K,t):=\exp\left(\sum_{m=1}^{\infty}\,\#\mathrm{Fix}^{*}(K,\phi^{\circ m})\cdot\frac{t^{m}}{m}\right),$
where $\mathrm{Fix}^{*}(K,\phi^{\circ m})$ represents the set of repelling fixed points of $\phi^{\circ m}$. It is a formal power series in $\mathbb{Z}[[t]]$, see §2.2.
Our first main result is concerned with identifying the conditions under which the Artin-Mazur zeta function $\zeta_{\phi}(K,t)$ is a rational function in $t$ over the integers $\mathbb{Z}$. To specify these conditions, let us denote the set of $K$-rational critical points of $\phi$ by $\mathrm{Crit}(\phi,K)$, that is, $\mathrm{Crit}(\phi,K)$ consists of the points $c\in\mathbb{P}^{1}(K)$ for which there is a sufficiently small neighborhood where $\phi$ is not bijective to a neighborhood of $\phi(c)$. Inspired by [5] and [10], we classify the map $\phi$ as hyperbolic if $\mathrm{Crit}(\phi,K)\cap\mathcal{J}(\phi,K)=\emptyset$, and we call $\phi$ subhyperbolic if all points in $\mathcal{J}(\phi,K)\cap\mathrm{Crit}(\phi,K)$ are eventually periodic under $\phi$. By definitions, any hyperbolic rational map is subhyperbolic.
Theorem 1.1.
Let $K$ be a complete non-archimedean field,and let $\phi\in K(z)$ be a subhyperbolic rational map of degree at least $2$ with nonempty and compact $\mathcal{J}(\phi,K)$. Then the Artin-Mazur zeta function $\zeta_{\phi}(K,t)$ is the quotient of two coprime polynomials of $t$ with coefficients in $\mathbb{Z}$.
If $\mathbb{P}^{1}(K)$ is compact, then $\mathcal{J}(\phi,K)$ is automatically compact. If $\mathbb{P}^{1}(K)$ is not compact (for instance, the residue field of $K$ is infinite), then $\mathcal{J}(\phi,K)$ could still be a compact subspace of $\mathbb{P}^{1}(K)$. We refer to [17, 19, 31] for examples of rational maps with compact Julia sets defined over a non-archimedean field with an infinite residue field.
We deduce Theorem 1.1 from a general framework for continuous maps on non-archimedean spaces, as discussed in §3. More precisely, we impose constrains on the singular points of such continuous maps (analogous to critical points of rational maps), and then establish the finiteness of periodic points for a given period, along with the rationality of the Artin-Mazur zeta function. A key ingredient is to encode the system $(\mathcal{J}(K,\phi),\phi)$ with a symbolic dynamical system $(\Sigma_{B},\sigma)$ of finite states, see §3.4. Additionally, we provide a precise formula for the Artin-Mazur zeta function in §3.5.
As a consequence of Theorem 1.1, the topological entropy $\mathfrak{h}(\phi,K)$ on the Julia set of a subhyperbolic map $\phi\in K(z)$ of degree at least $2$ with compact $\mathcal{J}(\phi,K)$ is shown to be the logarithm of an algebraic number. Furthermore, we demonstrate that this entropy is, in fact, the logarithm of a weak Perron number. Recall that a real positive algebraic integer $\lambda$ is called a weak Perron number if the absolute values of its Galois conjugates $\lambda^{\prime}\in\mathbb{R}$ satisfy $|\lambda^{\prime}|\leq\lambda$; equivalently, some power of $\lambda$ is a Perron number that is an algebraic integer $\geq 1$ whose remaining Galois conjugates have strictly smaller absolute value (see [21, Theorem 3] and [30, Proposition 5.1]). Then our result naturally raises the question: Which weak Perron numbers can be realized as the exponential of the topological entropy on the Julia set of a subhyperbolic map over a non-archimedean field?
Our subsequent result provides a realization of the set of weak Perron numbers as the topological entropy on the Julia sets for subhyperbolic maps in $K(z)$ with compact Julia sets, provided that $K$ is discretely valued, presenting an analogue to the result [30, Theorem 1.2] of Thurston.
Theorem 1.2.
Let $K$ be a complete non-archimedean field equipped with a discrete valuation. Then the set
$\left\{\exp(\mathfrak{h}(\phi,K)):\phi\in K(z)\text{ subhyperbolic with compact }\mathcal{J}(\phi,K)\subseteq\mathbb{P}^{1}(K)\right\}$
equals the set of weak Perron numbers.
The encoding of $(\mathcal{J}(\phi,K),\phi)$ with $(\Sigma_{B},\sigma)$ implies that the entropy $\mathfrak{h}(\phi,K)$ can be reformulated as
$\mathfrak{h}(\phi,K)=\limsup_{m\rightarrow\infty}\frac{1}{m}\log\#\mathrm{Fix}^{*}(K,\phi^{\circ m}).$
The Fatou components in $\mathcal{F}(K,\phi)$ are well-defined and classified in [6] though $\mathbb{P}^{1}(K)$ is totally disconnected. Since $\phi$ is eventually non-expending on any eventually periodic Fatou component, such a component carries zero entropy. Consequently, if $\mathbb{P}^{1}(K)$ has no wandering domains, then the entropy $\mathfrak{h}(\phi,K)$ equals the topological entropy $h_{\mathrm{top}}(\phi,K)$ of $\phi$ on $\mathbb{P}^{1}(K)$. In particular, if $\mathbb{P}^{1}(K)$ contains $22\deg\phi-$ 2 critical points of $\phi$, counted with multiplicities, then the subhyperbolicity of $\phi$ implies that $\mathcal{F}(K,\phi)$ has no wandering domain (see, e.g., [4, Theorem 1.2]), and hence $\mathfrak{h}(\phi,K)=h_{\mathrm{top}}(\phi,K)$; we refer to [7, Chapter 11] for a study of wandering domains. Moreover, the Poincaré recurrence theorem implies that any invariant probability measure on $\mathbb{P}^{1}(K)$ cannot charge the wandering domains, and consequently, provided that $\mathbb{P}^{1}(K)$ is compact, the variational principle yields $\mathfrak{h}(\phi,K)=h_{\mathrm{top}}(\phi,K)$.
Although our theorem 1.2 has a superficial resemblance to Thurston’s, our arguments differ from Thurston’s in key respects. In Thurston’s case, since $\mathbb{R}$ has a linear order, a Markov decomposition can be obtained directly from the forward orbit of the critical points, which induces that the exponential of any entropy is a weak Perron number. However, non-archimedean fields do not admit such linear orders, which makes the construction of a desired Markov decomposition is significantly more conceptual and complicated. Once such a Markov decomposition is obtained, as aforementioned, the direction in theorem 1.2 that $\exp(\mathfrak{h}(\phi,K))$ is a weak Perron number follows from the symbolic dynamics developed for Theorem 1.1. For the converse direction, given a weak Perron number $\lambda\geq 1$, we construct a natural (0,1)-matrix $A$ with leading eigenvalue $\lambda$ that exhibits properties reflecting the non-archimedean nature. We then define a piecewise linear map on finitely many disks in $K$ with $A$ as its transition matrix. By applying a result from [24], we appropriately glue the pieces of this map to obtain a rational map in $K(z)$. After verifying its dynamics, we conclude that this map is hyperbolic and has entropy $\log\lambda$. For more details, we refer to §4.2.
Outline. The paper is organized as follows. In §2, we provide the background material on non-archimedean subhyperbolic maps and the Artin-Mazur zeta function. In §3, assuming the compactness of the Julia set, we describe the dynamics on Julia set of a continuous map on a non-archimedean metric space under certain assumptions, named normality and homogeneousness, on the singular points. We begin by encoding the Julia dynamics with a symbolic dynamical system of countably many states and then upgrade it to a symbolic dynamical system of finitely many states. In this section, we demonstrate that the Artin-Mazur zeta function for such a continuous map is rational. In §4, we focus on the case of rational maps. We first prove Theorem 1.1 in §4.1 by verifying that the rational maps satisfy the assumptions in previous section. Then, in §4.2, we establish Theorem 1.2. Finally, in §5, we compute the Artin-Mazur zeta function and the entropy for a concrete example, using the method developed in this paper.
Acknowledgements. The work was partially done during the second author’s visit to the first author in the Department of Mathematics at National Taiwan Normal University (NTNU) and the third author’s visit to the second author in the Institute for Mathematical Sciences (IMS) at Stony Brook University. The authors are grateful to the Department of Mathematics at NTNU and the IMS for their generous hospitality. The third author is supported by Simons Collaboration Grant, Award Number 850685.
2. Preliminaries
In this section, we present the necessary background material. In §2.1, we introduce subhyperbolic continuous maps on non-archimedean spaces, extending the concept of subhyperbolic rational maps defined over non-archimedean fields. In §2.2, we discuss the Artin-Mazur zeta functions for continuous maps on non-archimedean spaces, for finite directed graphs and for subshifts of finite type, respectively.
2.1. Non-archimedean subhyperbolic maps
Let $X$ be a metric space associated with a metric $\rho$. We say that $X$ is non-archimedean if the metric $\rho$ satisfies the strong triangle inequality, that is, $\rho(x,y)\leq\max\{\rho(x,z),\rho(y,z)\}$ for any points $x,y,z\in X$. Given a non-archimedean metric space $(X,\rho)$, a closed disk in $X$ is a set of form
$\overline{D}(a,r):=\{x\in X:\rho(x,a)\leq r\}$
for some $a\in X$ and $r\geq 0$; and an open disk in $X$ is a set of form
$D(a,r):=\{x\in X:\rho(x,a)<r\}$
for some $a\in X$ and $r>0$. A disk is trivial if it is a singleton; and nontrivial if otherwise. We will use $U$ to denote a disk in $X$ without specifying whether it is open or closed, unless stated otherwise. Under the topology induced by $\rho$, any nontrivial disk in $X$ is a clopen set. Moreover, if two disks intersect, then one contains the other.
A continuous map $f:X\to X$ induces a non-archimedean dynamical system $(X,f)$. For any $n\geq 1$, write $f^{\circ n}$ for the $n$-times iterate of $f$. For a subset $Y$ of $X$, the grand orbit of $Y$ is
$\mathrm{GO}(Y):=\{x\in X:\exists\ n\geq 0,m\geq 0,\text{and}\ y\in Y\ \text{such that}\ f^{\circ n}(x)=f^{\circ m}(y)\}.$
A point $x\in X$ is a periodic point of period $n$ if $f^{\circ n}(x)=x$; in this case, its forward orbit $\{x,f(x),\dots,f^{\circ(n-1)}(x)\}$ is a periodic cycle of length $n$. If, in addition, $U$ is a proper subset of $f^{\circ n}(U)$ for any small disk $U$ containing $x$, we say that $x$ is a repelling periodic point and its forward orbit is a repelling periodic cycle. Given $n\geq 1$, denote by $\mathrm{Fix}^{\ast}(f^{\circ n})$ the set of all repelling periodic points of period $n$ for $f$. The Julia set $\mathcal{J}(f)$ of $f$ is the closure of the grand orbit of the repelling periodic points of $f$, that is,
$\mathcal{J}(f):=\overline{\mathrm{GO}(\cup_{n\geq 1}\mathrm{Fix}^{\ast}(f^{\circ n}))}.$
Then $f(\mathcal{J}(f))$ is subset of $\mathcal{J}(f)$, and hence one obtains the subsystem $(J(f),f)$. We mention that $\mathcal{J}(f)$ could be empty.
Given a disk $U$ in $X$, we say that $f$ is a (distance) scaling map on $U$ if the following hold:
  1. (1)
    $f(U)$ is a disk in $X$, and
  2. (2)
    there exists $\lambda>0$ such that for any $x,y\in U$,
    $\rho(f(x),f(y))=\lambda\cdot\rho(x,y).$
In this case, we call $U$ a (distance) scaling disk for $f$ and observe that $f$ is a bijection from $U$ onto $f(U)$. Observe that any trivial disk is a scaling disk.
A point $x\in\mathcal{J}(f)$ is regular if there exists a disk containing $x$ where $f$ is a scaling map. Otherwise, $x\in\mathcal{J}(f)$ is singular. Denote by $\mathrm{Sing}(f)$ the set of singular points of $f$ in $\mathcal{J}(f)$.
Definition 2.1.
Given a non-archimedean dynamical system $(X,f)$ with non-empty $\mathcal{J}(f)$, we say that
  1. (1)
    $f$ is hyperbolic if $\mathrm{Sing}(f)$ is empty;
  2. (2)
    $f$ is subhyperbolic if $\mathrm{Sing}(f)$ is finite and, for any $x\in\mathrm{Sing}(f)$, there is $k_{x}>0$ such that $f^{\circ k_{x}}(x)$ is a repelling periodic point of $f$.
Trivially, if $f$ is hyperbolic, then it is subhyperbolic. Moreover, if $f$ is a subhyperbolic rational map with compact $\mathcal{J}(f)$, then $J(f)$ admits some expanding (singular) metric, see [11].
In the context of rational maps $\phi\in K(z)$, recall from §1 the set of critical points $\mathrm{Crit}(\phi,K)$ and the Julia set $\mathcal{J}(\phi,K)$. Regarding $\phi$ as a continuous map defined on the non-archimedean space $\mathbb{P}^{1}(K)$, we immediately have that
$\mathrm{Crit}(\phi,K)\cap\mathcal{J}(\phi)=\mathrm{Sing}(\phi).$
Moreover, if $\phi$ is subhyperbolic, then $\mathcal{J}(\phi,K)$ coincides to $\mathcal{\mathcal{J}(\phi)}$:
Lemma 2.2.
Let $\phi\in K(z)$ be a subhyperbolic rational map of degree at least $2$. Then
$\mathcal{J}(\phi,K)=\mathcal{\mathcal{J}(\phi)}.$
In particular,
$\mathrm{Crit}(\phi,K)\cap\mathcal{J}(\phi,K)=\mathrm{Sing}(\phi).$
Proof.
Since $\mathcal{F}(\phi,K)$ is forward invariant under $\phi$ by definition, it follows that $\mathcal{J}(\phi,K)$ is backward invariant under $\phi$. Noting that $\mathrm{Fix}^{\ast}(\phi^{\circ n})$ is a subset of $\mathcal{J}(\phi,K)$ for any $n\geq 1$, since $\mathcal{J}(\phi,K)$ is closed, we conclude that $\mathcal{\mathcal{J}(\phi)}\subseteq\mathcal{J}(\phi,K)$.
Now pick $x\in\mathcal{J}(\phi,K)$ and consider a sufficiently small neighborhood $U$ of $x$ in $\mathbb{P}^{1}(K)$. By the non-archimedean Montel’s Theorem, see [12, Corollary D] (see also [18]), there exists $n:=n_{U}\geq 1$ such that $U$ is a subset of $\phi^{\circ n_{U}}(U)$. If $\phi^{\circ n}$ is a scaling map on $U$, then $U$ contains a point in $\mathrm{Fix}^{\ast}(\phi^{\circ n})$. If $\phi^{\circ n}$ is not a scaling map on $U$ for any neighborhood $U$ of $x$, then $x$ is an accumulation point of the iterated preimages of a critical point, say $c$, in $\mathrm{Crit}(\phi,K)\cap\mathcal{J}(\phi,K)$. Since $\phi$ is subhyperbolic, the point $c$ is contained in $\mathrm{GO}(\cup_{n\geq 1}\mathrm{Fix}^{\ast}(\phi^{\circ n}))$. It follows that $x\in\mathcal{\mathcal{J}(\phi)}$. Hence $\mathcal{J}(\phi,K)\subseteq\mathcal{\mathcal{J}(\phi)}$. ∎
2.2. Artin-Mazur zeta function
In this section, we state the Artin-Mazur zeta function in several settings. We refer to [26, 27] for references.
For a continuous map $f:X\to X$ on a non-archimedean space $X$, if $\mathrm{Fix}^{*}(f^{\circ m})$ is a finite set for each $m\geq 1$, the Artin-Mazur zeta function on $\mathcal{J}(f)$ is defined by
$\zeta_{f}(X,t):=\exp\left(\sum_{m=1}^{\infty}\,\#\mathrm{Fix}^{*}(f^{\circ m})\cdot\frac{t^{m}}{m}\right).$
In fact, $\zeta_{f}(X,t)\in\mathbb{Z}[[t]]$ from the product formula
$\zeta_{f}(X,t)=\prod_{\text{repelling cycles}\ \gamma}(1-t^{\#\gamma})^{-1}.$
The Artin-Mazur zeta function is also well-defined for finite directed graphs. Consider a finite directed graph $G$. Let $A_{G}$ be its transition matrix of $G$, that is, each entry of $A_{G}$ represents the number of directed edges of $G$ connecting the corresponding vertices. Then $A_{G}$ has non-negative integer entries. The Artin-Mazur zeta function for $G$ is defined by
$\zeta(G,t):=\exp\left(\sum_{m=1}^{\infty}\,\mathcal{N}_{m}(G)\cdot\frac{t^{m}}{m}\right),$
where $\mathcal{N}_{m}(G)$ is the number of loops consisting of $m$ (not necessarily distinct) directed edges. It follows that $\mathcal{N}_{m}(G)=trace((A_{G})^{m})$ and
$\zeta(G,t)=\det(I-tA_{G})^{-1},$
where $I$ is the identity matrix with same size of $A_{G}$. Let $\lambda\geq 1$ be the leading eigenvalue of $A_{G}$ (in fact, $\lambda$ is a weak Perron number). Then, as a function in $\mathbb{C}$, the function $\zeta(G,t)$ is non-zero and analytic for $|t|<1/\lambda$ and has a simple pole at $t=1/\lambda$.
Moreover, given a finite matrix $A$ with non-negative integer entries, its transition graph $G_{A}$ is a finite directed graph such that $A_{G}=A$. Then
$\zeta(G_{A},t)=\det(I-tA)^{-1}.$
Now for symbolic dynamical systems, consider a subshift of finite type $(\Sigma_{B},\sigma)$ with transition matrix $B$, where $\sigma$ is the left shift on $\Sigma_{B}$. The Artin-Mazur zeta function on $\Sigma_{A}$ is defined by
$\zeta_{\sigma}(\Sigma_{B},t):=\exp\left(\sum_{m=1}^{\infty}\,\#\mathrm{Fix}_{B}(\sigma^{\circ m})\cdot\frac{t^{m}}{m}\right),$
where $\mathrm{Fix}_{B}(\sigma^{\circ m})$ is the set of the periodic points of $\sigma$ of period $m$ in $\Sigma_{B}$. Then
$\zeta_{\sigma}(\Sigma_{B},t)=\zeta(G_{B},t)=\det(I-tB)^{-1}.$
Now consider a continuous map $f:X\to X$ on a non-archimedean space $X$ and a subshift of finite type $(\Sigma_{B},\sigma)$. If $(\mathcal{J}(f),f)$ is topologically conjugate to $(\Sigma_{B},\sigma)$, then
$\zeta_{f}(X,t)=\zeta_{\sigma}(\Sigma_{B},t)=\det(I-tB)^{-1};$
moreover, the topological entropy $\mathfrak{h}(f)$ for $(\mathcal{J}(f),f)$ and the topological entropy $\mathfrak{h}_{B}(\sigma)$ for $(\Sigma_{B},\sigma)$ satisfy that
$\mathfrak{h}(f)=\mathfrak{h}_{B}(\sigma)=\log\lambda_{B}=\limsup_{m\rightarrow\infty}\frac{1}{m}\log\#\mathrm{Fix}_{B}(\sigma^{\circ m})=\limsup_{m\rightarrow\infty}\frac{1}{m}\log\#\mathrm{Fix}^{*}(f^{\circ m}),$
where $\lambda_{B}\geq 1$ is the leading eigenvalue of $A$.
3. Rationality of the zeta function
In this section, we always assume that $(X,\rho)$ is a complete non-archimedean metric space and $f:X\to X$ is a subhyperbolic continuous map with non-empty Julia set $\mathcal{J}(f)$. For each periodic cycle $\gamma$ of $f$ in $X$, recall that $\mathrm{GO}(\gamma)$ is the grand orbit of $\gamma$; equivalently, the backward orbit of $\gamma$. Recall the set of singular points $\mathrm{Sing}(f)$ contained in $\mathcal{J}(f)$ and denote by $\Gamma(f)$ the set of all repelling cycles $\gamma$ of $f$ such that $\mathrm{GO}(\gamma)$ intersects $\mathrm{Sing}(f)$. Then $\Gamma(f)$ is a finite set since $\mathrm{Sing}(f)$ is finite. Moreover, for each $\gamma\in\Gamma(f)$, we mark a point $p_{\gamma}\in\gamma$. Our goal is to establish the rationality of the zeta function for $f$ under some assumptions on $\mathcal{J}(f)$ and $\mathrm{Sing}(f)$; see §3.1 for the precise statement.
3.1. Definitions and main result
First, we introduce some constrains on the behavior of the singular points of $f$.
Definition 3.1.
A point $x\in\mathrm{Sing}(f)$ is normal if there exist a disk $U\subseteq X$ of $x$ and a positive number $r_{x}>0$ such that
  1. (1)
    $x$ is the unique singular point in $U$, that is, $U\cap\mathrm{Sing}(f)=\{x\}$,
  2. (2)
    $f(y)\not=f(x)$ for any $y\in U\setminus\{x\}$,
  3. (3)
    any maximal scaling disk $V$ in $U\setminus\{x\}$ has diameter at least $r_{x}\cdot\rho(x,V)$, and
  4. (4)
    the image $f(V)$ of any maximal scaling disk $V$ in $U\setminus\{x\}$ has diameter at least $r_{x}\cdot\rho(f(x),f(V))$.
Moreover, such a disk $U$ is called a normal disk for the singular point $x$.
For $x\in\mathrm{Sing}(f)$, let $\gamma:=\gamma_{x}\in\Gamma(f)$ such that $x\in\mathrm{GO}(\gamma)$ and denote by $\ell_{x}\geq 1$ the smallest integer such that $f^{\circ\ell_{x}}(x)=p_{\gamma}$.
Definition 3.2.
We say that $f$ satisfies the homogeneous condition if there exists a map $\varphi:\Gamma(f)\to\mathbb{Z}_{>0}$ and, for each $z\in\mathrm{Sing}(f)$, there exist a disk $V_{z}\subset X$ containing $z$ and a scaling map $T_{z}:V_{z}\to X$ fixing $z$ such that the following hold: given any $x\in\mathrm{Sing}(f)$,
  1. (1)
    in $V_{x}$,
    $f^{\circ\ell_{x}}\circ T_{x}=f^{\circ(\varphi(\gamma_{x})+\ell_{x})};$
  2. (2)
    if $y:=f^{\circ j}(x)\in\mathrm{Sing}(f)$ for some $j\geq 1$, then $f^{\circ j}(V_{x})\subseteq V_{y}$ and
    $f^{\circ j}\circ T_{x}=T_{y}\circ f^{\circ j}.$
See Figure 1 for an illustration of Definition 3.2.
Figure 1. The map $T_{x}$ in Definition 3.2.
If $f$ satisfies the homogeneous condition, then the appropriate $\varphi(\gamma)$ is a multiple of the length of $\gamma\in\Gamma(f)$. The homogeneous condition restricts the distribution of the Julia set $\mathcal{J}(f)$ near the singular points. More precisely, given an annulus $A_{x}\subseteq U_{x}$ around a singular point $x$, the intersection $A_{x}\cap\mathcal{J}(f)$ is transported by $T_{x}$ in $V_{x}$, that is,
$T_{x}(T_{x}^{-1}(A_{x})\cap\mathcal{J}(f))=A_{x}\cap\mathcal{J}(f).$
Moreover, for a small disk $V\subseteq V_{x}\setminus\{x\}$ such that $f^{\circ\ell_{x}}$ is scaling in $V$ and in $T_{x}(V)$, then
$T_{x}|_{V}=(f^{\circ\ell_{x}}|_{T_{x}(V)})^{-1}\circ f^{\circ(\ell_{x}+\varphi(\gamma_{x}))}.$
Furthermore, the homogeneous condition can be extended to the forward orbits of the singular points, see §3.4.1.
Remark 3.3.
If $f$ is a subhyperbolic rational map defined over a non-archimedean field, then each of its singular points (the critical points in $\mathcal{J}(f)$) is normal and $f$ satisfies the homogeneous condition, see §4.1.
The main result in this section is the following rationality result; its proof relies on a series of results given in next subsections.
Theorem 3.4.
Let $f:X\to X$ be a non-archimedean subhyperbolic continuous map. Assume that
  1. (1)
    $\mathcal{J}(f)$ is compact,
  2. (2)
    each point in $\mathrm{Sing}(f)$ is normal, and
  3. (3)
    $f$ satisfies the homogeneous condition.
Then for each $n\geq 1$,
#$\#\mathrm{Fix}^{*}(f^{\circ n})<+\infty.$
Moreover, $\zeta_{f}(X,t)=P(t)/Q(t)$ for some coprime polynomials $P,Q$ in $\mathbb{Z}[t]$.
Under the assumptions in Theorem 3.4, we will describe the dynamics of $f$ on $J(f)$ with a finite Markov partition, which in fact leads an exact formula for $\zeta_{f}(X,t)$, see §3.4 and §3.5.
In following subsections, we aim to establish Theorem 3.4. We first analyze the dynamics along the orbits of singular points in §3.2. We then obtain a natural countable Markov partition of $\mathcal{J}(f)$ in §3.3 using a process called splitting. In §3.4, we upgrade the countable Markov partition to a finite Markov partition using a factor map. We finally conclude Theorem 3.4 in §3.5.
3.2. Dynamics along the orbits of singular points
In this subsection, we assume that $\mathcal{J}(f)$ is compact and each point in $\mathrm{Sing}(f)$ is normal. For $\gamma\in\Gamma(f)$, define
$\mathrm{GO}^{*}(\gamma):=\{f^{\circ j}(x):x\in\mathrm{GO}(\gamma)\cap\mathrm{Sing}(f),j\geq 0\}.$
Obviously $\mathrm{GO}^{*}(\gamma)\subseteq\mathrm{GO}(\gamma)$. Under the compactness of $\mathcal{J}(f)$ and the normality of $\mathrm{Sing}(f)$, we construct a neighborhood of $\mathrm{GO}^{*}(\gamma)$ so that, off $\mathrm{GO}^{*}(\gamma)$, it contains countably many scaling disks intersecting $\mathcal{J}(f)$:
Lemma 3.5.
There exist a neighborhood $\overline{W}(\gamma)$ of $\mathrm{GO}^{*}(\gamma)$, consisting of finitely many closed disks, in $X$ and a collection $\mathcal{U}(\gamma)$ of countably many scaling disks such that the following hold:
  1. (1)
    For each $U\in\mathcal{U}(\gamma)$, $U\cap\mathcal{J}(f)\not=\emptyset$ and $U\subseteq\overline{W}(\gamma)$.
  2. (2)
    $\bigcup_{U\in\mathcal{U}(\gamma)}U$ covers $\left(\overline{W}(\gamma)\cap\mathcal{J}(f)\right)\setminus\mathrm{GO}^{*}(\gamma)$.
  3. (3)
    For any $U\in\mathcal{U}(\gamma)$, either $f(U)\in\mathcal{U}(\gamma)$ or $f(U)\cap\overline{W}(\gamma)=\emptyset$.
Proof.
Pick a closed disk $\overline{D}$ of the marked point $p:=p_{\gamma}$ in $X$ and pull back $\overline{D}$ along the orbits in $\mathrm{GO}^{*}(\gamma)$. For each point $x\in\mathrm{GO}^{*}(\gamma)$, we obtain a disk $\overline{D}_{x}$, that is, if $f^{\circ\ell}(x)=p$ for a smallest integer $\ell:=\ell_{x}\geq 1$, then $\overline{D}_{x}$ is the large disk around $x$ such that $f^{\circ\ell}(\overline{D}_{x})\subseteq\overline{D}$. Denoting by $s\geq 1$ the length of $\gamma$, we can shrink $\overline{D}$ so that the following hold:
  1. (1)
    The disk $\overline{D}=\overline{D}_{p}$ is a scaling disk of $f^{\circ s}$.
  2. (2)
    The disks $\overline{D}_{x},x\in\mathrm{GO}^{*}(\gamma)$ are pairwise disjoint.
  3. (3)
    For any $y\in\gamma$, if $f^{\circ s}(\overline{D}_{y})\cap\overline{D}_{x}\not=\emptyset$, then $x=y$.
  4. (4)
    If $\overline{D}_{x}\cap\mathrm{Sing}(f)=\emptyset$, then $\overline{D}_{x}$ is a scaling disk of $f$.
  5. (5)
    If $\overline{D}_{x}\cap\mathrm{Sing}(f)\not=\emptyset$ , then $x$ is the unique singular point in $\overline{D}_{x}$ and $\overline{D}_{x}$ is a normal disk of $x$ (see Definition 3.1).
Set
$\overline{W}(\gamma):=\bigcup_{x\in\mathrm{GO}^{*}(\gamma)}\overline{D}_{x},$
and observe that for any $y\in\gamma$, the disk $\overline{D}_{y}$ is a scaling disk of $f^{\circ s}$.
For each $x\in\mathrm{GO}^{*}(\gamma)\cap\mathrm{Sing}(f)$, recall from Definition 3.1 the number $r_{x}>0$. Fix
$0<\delta<\prod_{x\in\mathrm{GO}^{*}(\gamma)\cap\mathrm{Sing}(f)}r_{x}.$
In $\overline{D}_{p}=\overline{D}$, we consider
$\mathcal{U}_{p}:=\left\{\text{maximal closed disks}\ U\subseteq\overline{D}_{p}\setminus\{p\}:\mathrm{diam}(U)\leq\delta\cdot\rho(p,U)\ \text{and}\ U\cap\mathcal{J}(f)\not=\emptyset\right\}.$
For any $x\in\mathrm{GO}^{*}(\gamma)$, let $U\subseteq\overline{D}_{x}\setminus\{x\}$ be the maximal closed disk such that $f^{\circ\ell_{x}}(U)$ is contained in a disk $V\in\mathcal{U}_{p}$. By the normality of the singular points, we claim that $f^{\circ\ell_{x}}$ is a scaling in $U$ and $f^{\circ\ell_{x}}(U)=V\in\mathcal{U}_{p}$. Indeed, since $\mathrm{diam}(V)\leq\delta\cdot\rho(p,V)$, in view of Definition 3.1 (4), there exists a maximal scaling disk $V^{\prime}\subseteq\overline{D}_{x}$ such that $V^{\prime}\cap U\not=\emptyset$ and $V\subseteq f^{\circ\ell_{x}}(V^{\prime})$, which implies $U\subseteq V^{\prime}$ and hence validates the claim. We now set
$\mathcal{U}_{x}:=\left\{\text{closed disks}\ U\subseteq\overline{D}_{x}\setminus\{x\}:f^{\circ n_{x}}(U)\in\mathcal{U}_{p}\ \text{and}\ U\cap\mathcal{J}(f)\not=\emptyset\right\}.$
Let
$\mathcal{U}(\gamma):=\bigcup_{x\in\mathrm{GO}^{*}(\gamma)}\mathcal{U}_{x}.$
Then for any $U\in\mathcal{U}(\gamma)$, we have $U\subseteq\overline{W}(\gamma)$ and $U\cap\mathcal{J}(f)\not=\emptyset$. Moreover, $\mathcal{U}(f)$ forms a covering of $\left(\bigcup_{x\in\mathrm{GO}^{*}(\gamma)}\overline{D}_{x}\right)\setminus\mathrm{GO}^{*}(\gamma)$. Thus statements (1) and (2) hold.
Pick any $U\in\mathcal{U}(\gamma)$. If $U$ is not a subset of $\overline{D}_{p}\setminus f^{-s}(\overline{D}_{p})$, then $f(U)$ is contained in $\mathcal{U}(\gamma)$. If $U$ is a subset of $\overline{D}_{p}\setminus f^{-s}(\overline{D}_{p})$, then $f(U)$ is a disk contained in $f(\overline{D}_{p})\setminus\overline{D}_{f(p)}=f^{\circ s}(\overline{D}_{f(p)})\setminus\overline{D}_{f(p)}$, and hence $f(U)$ is not contained in $\overline{D}(\gamma)$. Thus statement (3) holds.
To complete the proof, we must show the collection $\mathcal{U}(\gamma)$ is countable. It is sufficient to show that $\mathcal{U}_{x}$ is countable for each $x\in\mathrm{GO}^{*}(\gamma)$. For each large $n\geq 1$, consider the open disk $D_{x,n}:=D(x,1/n)$. Pick a sufficiently large $n_{0}\gg 1$ so that $\overline{D}_{x,n}\subsetneq\overline{D}_{x}$, and consider the (closed) annuli $\overline{A}_{x,n_{0}}:=\overline{D}_{x}\setminus D_{x,n_{0}}$ and $\overline{A}_{x,n+1}=\overline{D}_{x,n}\setminus D_{x,n+1}$ for $n\geq n_{0}$. Then for each $n\geq n_{0}$, the intersection $\overline{A}_{x,n}\cap\mathcal{J}(f)$ is compact, and hence $\mathcal{U}_{x}$ contains finitely many elements which are subsets of $\overline{A}_{x,n}$. Since the union of the annuli $\overline{A}_{x,n},n\geq n_{0}$ is $\overline{D}_{x}\setminus\{x\}$, we conclude that $\mathcal{U}_{x}$ is countable. ∎
In next subsection, we will take into account the part of $\mathcal{J}(f)$ not covered by $\cup_{\gamma\in\Gamma(f)}\mathcal{U}(\gamma)$ and obtain a covering of $\mathcal{J}(f)$ satisfying the Markov property. Here, we say a collection $\mathcal{U}$ of disjoint disks in $X$ satisfies the Markov property (with respect to $f$), if $f(U)$ contains a disk in $\mathcal{U}$ for any $U\in\mathcal{U}$. This covering will induce a symbolic dynamical system to record the dynamics of $f$ in $\mathcal{J}(f)$
3.3. Countable symbols via splitting
In this subsection, we continue to assume that $\mathcal{J}(f)$ is compact and each point in $\mathrm{Sing}(f)$ is normal. Inspired by [10], we construct a symbolic dynamical system of countably many states, applying a process so-called splitting, and establish a topological conjugacy result:
Proposition 3.6.
There exist a countable set $\mathcal{U}_{\ast}$ consisting of scaling disks of $f$, a symbolic dynamical system $(\Sigma_{A},\sigma)$ of countably many states corresponding to the disks in $\mathcal{U}_{\ast}$, and a bijection $h:\mathcal{J}(f)\to\Sigma_{A}$ such that $h$ is a topological conjugation in $\mathcal{J}(f)\setminus\left(\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)\right)$ between $f$ and $\sigma$
The main ingredient to prove Proposition 3.6 is to construct a collection of disjoint disks with the Markov property from a given collection of disks. The main technique we apply is the splitting:
Definition 3.7.
  1. (1)
    Given a set $Y\subseteq X$, for nontrivial disks $U\subseteq X$ and $U_{0}\subsetneq U$ intersecting $Y$, we say that the collection
    $\left\{U_{0},V:V\ \text{is a maximal disk in}\ U\backslash U_{0}\ \text{intersecting}\ Y\right\}$
    is the splitting of $U$ via $U_{0}$ with respect to $Y$.
  2. (2)
    Given a set $Y\subseteq X$, for a collection $\mathcal{U}_{0}$ of disjoint disks, each intersecting $Y$, in $X$, we say that a collection $\mathcal{U}$ of disjoint disks is a splitting of $\mathcal{U}_{0}$ if $\mathcal{U}$ can be obtained by replacing a single nontrivial disk in $\mathcal{U}_{0}$ with its splitting via a nontrivial proper subdisk with respect to $Y$. Moreover, for an integer $n\geq 1$, we say that a collection $\mathcal{U}^{\prime}$ of disjoint disks is an n-step splitting of $\mathcal{U}_{0}$ if there exists a sequence of collection $\{\mathcal{U}_{i}\}_{i=0}^{n}$ of disjoint disks such that $\mathcal{U}_{i+1}$ is a splitting of $\mathcal{U}_{i}$ for $10\leq i\leq n-$ 1 and $\mathcal{U}^{\prime}=\mathcal{U}_{n}$.
Given a collection of disjoint scaling disks such that all but finitely many disks have images containing disks in the collection, in next result (Lemma 3.8), we show that this collection can be upgraded by a finite-step splitting to a collection of disjoint scaling disks with the Markov property. To achieve this, let us first introduce an index for a general continuous map $\psi:X\to X$ with nonempty $\mathcal{J}(\psi)$.
Let $\mathcal{U}$ be a collection of disjoint scaling disks of $\psi$ covering $\mathcal{J}(\psi)$ such that at least one disk is nontrivial and each disk intersects $\mathcal{J}(\psi)$. For any nontrivial disk $U\in\mathcal{U}$, considering a forward iterate of $U$ if necessary, we can assume that $U$ contains a repelling periodic point, say $p$, of period $s$. Then either there exists a smallest integer $11\leq m\leq s-$ 1 such that $\psi^{\circ m}(U)$ contains a disk in $\mathcal{U}$ or $\psi^{\circ i}(U)$ is properly contained in a disk in $\mathcal{U}$ for each $11\leq i\leq s-$ 1. In the latter case, we have that $\psi^{\circ s}$ is a scaling in $U$ and $p\in U\subsetneq\psi^{\circ s}(U)$, which implies that $\psi^{\circ s}(U)\setminus U$ contains an element in $\mathcal{U}$ as a subset. Thus for each nontrivial disk $U\in\mathcal{U}$, there exists a smallest integer $m:=m_{\psi}(U)\geq 1$ such that $\psi^{\circ m}(U)$ contains a disk in $\mathcal{U}$.
We define the index
$M_{\psi}(\mathcal{U}):=\sum_{\text{nontrivial}\ U\in\mathcal{U}}(m_{\psi}(U)-1).$
If there exists a finite subcollection $\mathcal{U}^{\prime}$ of $\mathcal{U}$ consisting of nontrivial disks such that $\psi(U)$ contains a disk in $\mathcal{U}$ for any $U\in\mathcal{U}\setminus\mathcal{U}^{\prime}$, then
$0\leq M_{\psi}(\mathcal{U})<+\infty.$
This index indicates the number of splittings for $\mathcal{U}$ to produce a collection satisfying the Markov property:
Lemma 3.8.
Consider a continuous map $\psi:X\to X$ with nonempty $\mathcal{J}(\psi)$ and let $\mathcal{U}$ be a collection of disjoint scaling disks, each intersecting $\mathcal{J}(\psi)$, of $\psi$ covering $\mathcal{J}(\psi)$. Assume that there exists a finite subcollection $\mathcal{U}^{\prime}$ of $\mathcal{U}$ consisting of nontrivial disks such that $\psi(U)$ contains a disk in $\mathcal{U}$ for any $U\in\mathcal{U}\setminus\mathcal{U}^{\prime}$. Then there exists an $M_{\psi}(\mathcal{U})$-step splitting of $\mathcal{U}$ satisfying the Markov property.
Proof.
We can assume that $\psi(U)$ does not contain a disk in $\mathcal{U}_{0}:=\mathcal{U}$ for any $U\in\mathcal{U}^{\prime}$; so $\phi(U)$ is contained in some disk in $\mathcal{U}_{0}$ since $\mathcal{J}(\phi)$ is forward invariant. Then the index
$0<M_{\psi}({\mathcal{U}_{0}})<+\infty.$
Consider $U\in\mathcal{U}^{\prime}$ and the smallest integer $m:=m(U)$ such that $\psi^{\circ m}(U)$ contains a disk in $\mathcal{U}_{0}$. Let $U^{\prime}\in\mathcal{U}_{0}$ such that $\psi^{\circ(m-1)}(U)\subsetneq U^{\prime}$. The splitting of $U^{\prime}$ via $\psi^{\circ(m-1)}(U)$ with respect to $\mathcal{J}(\phi)$ induces a splitting of $\mathcal{U}_{0}$, denoted by $\mathcal{U}_{1}$. Observe that $\psi^{\circ(m-1)}(U)\in\mathcal{U}_{1}$. We claim
$M_{\psi}(\mathcal{U}_{1})=M_{\psi}(\mathcal{U}_{0})-1.$
It suffices to show that $\psi(V)$ contains a disk in $\mathcal{U}_{1}$ for any maximal disk $V\subseteq U^{\prime}\setminus\psi^{\circ(m-1)}(U)$. Suppose otherwise. Then there exist a disk in $\mathcal{U}_{1}$ and hence a disk $V^{\prime}\in\mathcal{U}_{0}$ such that $\psi(V)\subsetneq V^{\prime}$. Observe that $V^{\prime}\subseteq\psi(U^{\prime})$ since $V\subseteq U^{\prime}$ and $\psi^{\circ m}(U)\subseteq\psi(U^{\prime})$ contains a disk in $\mathcal{U}_{0}$. Noting that $\psi$ is a scaling in $U^{\prime}$, we conclude that $(\psi|_{U^{\prime}})^{-1}(V^{\prime})$ is contained in a maximal disk in $U^{\prime}\setminus\psi^{\circ(m-1)}(U)$ and hence $(\psi|_{U^{\prime}})^{-1}(V^{\prime})\subseteq V$, which implies that $V^{\prime}\subseteq\phi(V)$. This is a contradiction.
Repeatedly applying the splitting process, since $M_{\psi}(\mathcal{U}_{0})$ is finite, we obtain a finite-step splitting $\mathcal{U}_{n}$ of $\mathcal{U}_{0}$ such that $M_{\psi}({\mathcal{U}_{n}})=0$, where $n=M_{\psi}(\mathcal{U}_{0})$. Then $\mathcal{U}_{n}$ satisfies the Markov property. ∎
We proceed to prove Proposition 3.6.
Proof of Proposition 3.6.
Write $\Gamma(f):=\{\gamma_{1},\dots,\gamma_{i_{0}}\}$ and consider the collections $\mathcal{U}(\gamma_{i}),1\leq i\leq i_{0}$ as in Lemma 3.5. Set
$\mathcal{U}:=\bigcup_{i=1}^{i_{0}}\left(\mathcal{U}(\gamma_{i})\cup\left\{\{x\}:x\in\mathrm{GO}^{*}(\gamma_{i})\right\}\right).$
Then $\mathcal{U}$ the collection of disjoint scaling disks of $f$. Since $\mathcal{J}(f)$ is compact, there exists a collection $\mathcal{U}^{\prime}$ of finitely many nontrivial scaling disks, each intersecting $\mathcal{J}(f)$, of $f$ covering the complement $\mathcal{J}(f)\setminus(\bigcup_{U\in\mathcal{U}}U)$. Then the collection
$\mathcal{U}_{0}:=\mathcal{U}\cup\mathcal{U}^{\prime}$
covers $\mathcal{J}(f)$ and each disk in $\mathcal{U}_{0}$ is a scaling disk of $f$. Moreover, all disks but finitely many nontrivial disks in $\mathcal{U}_{0}$ have images containing a disk in $\mathcal{U}_{0}$. Thus by Lemma 3.8, there exists a collection $\mathcal{U}_{\ast}$ of disjoint scaling disks, each intersecting $\mathcal{J}(f)$, of $f$ covering $\mathcal{J}(f)$ and satisfying the Markov property. Moreover, since $\mathcal{J}(f)$ is compact, we conclude that $\mathcal{U}_{\ast}$ contains countably many disks.
Regard the disks in $\mathcal{U}_{\ast}$ as symbols, and let $A$ be the transition matrix for the disks in $\mathcal{U}_{\ast}$. Then we obtain a symbolic dynamical system $(\Sigma_{A},\sigma)$. Define the map
$h:\mathcal{J}(f)\to\Sigma_{A}$
sending $x\in\mathcal{J}(f)$ to $(\alpha_{0},\alpha_{1},\alpha_{2},\dots)\in\Sigma_{A}$, where $\alpha_{i}$ corresponds to the disk in $\mathcal{U}_{\ast}$ containing $f^{\circ i}(x)$ for each $i\geq 0$.
We now show that $h$ is a injection. Suppose that $h(x)=h(y)$ for $x\not=y\in\mathcal{J}(f)$. Then both $x$ and $y$ are not in the forward orbits of the singular points. Let $\overline{D}$ be the smallest closed disk containing both $x$ and $y$ in $X$. Then $\overline{D}\cap\mathcal{J}(f)\not=\emptyset$. Hence considering a forward iterate of $\overline{D}$ if necessary, we can assume that $\overline{D}$ contains a repelling periodic point, say $p$, of period $s\geq 1$. Observe that $f^{\circ i}(\overline{D})$ is contained in a scaling disk in $\mathcal{U}_{\ast}$ for each $i\geq 0$ since $h(x)=h(y)$. We conclude that $\overline{D}\subsetneq f^{\circ s}(\overline{D})$. It follows that there exists $i_{0}\geq 1$ such that $f^{\circ(i_{0}s)}(\overline{D})$ contains the disk in $\mathcal{U}_{\ast}$ containing $\overline{D}$. Thus the $(i_{0}s)$-th entries of $h(x)$ and $h(y)$ are distinct, which contradicts that $h(x)=h(y)$. Therefore, $h$ is injective.
We now show that $h$ is a surjection. Pick $\alpha:=(\alpha_{0},\alpha_{1},\dots)\in\Sigma_{A}$. If some $\alpha_{j}$ corresponds to the singleton of a point in the forward orbit of $\mathrm{Sing}(f)$, then the scaling property of $f$ in the disks corresponding to $\alpha_{i},0\leq i\leq j-1$ implies that there exists $x\in\mathcal{J}(f)$ such that $h(x)=\alpha$. We now assume that each $\alpha_{i}$ corresponding to a nontrivial disk $U_{i}\in\mathcal{U}_{\ast}$. Observe that there exists $\ell_{i}\geq 1$ such that $f^{\circ\ell_{i}}(U_{i})$ contains at least two disks in $\mathcal{U}_{\ast}$; indeed, for otherwise, considering a forward iterate of $U_{i}$ if necessary, we can assume that $U_{i}$ contains a repelling periodic point and conclude that $U_{i}\subsetneq f^{\circ s_{i}}(U_{i})$ for some $s_{i}\geq 1$, which contradicts that each forward image of $U_{i}$ is a disk in $\mathcal{U}_{\ast}$. Now set $m_{0}=0$ and $m_{i}=\ell_{m_{i-1}}$ for each $i\geq 1$. Considering the map
$f^{-m_{i}}:U_{\sum_{j=0}^{i}m_{j}}\cap\mathcal{J}(f)\to U_{\sum_{j=0}^{i-1}m_{j}}\cap\mathcal{J}(f)$
for each $i\geq 1$, we have that
$f^{-m_{i}}(U_{\sum_{j=0}^{i}m_{j}}\cap\mathcal{J}(f))\subsetneq U_{\sum_{j=0}^{i-1}m_{j}}\cap\mathcal{J}(f)$
and conclude that the preimage
$f^{-\sum_{j=1}^{i}m_{j}}(U_{\sum_{j=0}^{i}m_{j}}\cap\mathcal{J}(f))=\left(\prod_{j=1}^{i}f^{-m_{j}}\right)\left(U_{\sum_{j=0}^{i}m_{j}}\cap\mathcal{J}(f)\right)$
shrinks to a point, say $x$, in $U_{0}\cap\mathcal{J}(f)$, as $i\to\infty$, since $\mathcal{J}(f)$ is compact. It follows that $h(x)=\alpha$.
We claim that $h\circ f=\sigma\circ h$ in $\mathcal{J}(f)$. Indeed, for $x\in\mathcal{J}(f)$, writing $h(x)=(\alpha_{0},\alpha_{1},\alpha_{2},\dots)\in\Sigma_{A}$, we obtain that
$h\circ f(x)=(\alpha_{1},\alpha_{2},\dots)=\sigma\circ h(x).$
Thus, to complete the proof, it suffices to show that $h$ is continuous in $\mathcal{J}(f)\setminus\left(\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)\right)$ and $h^{-1}$ is continuous in $\Sigma_{A}\setminus h\left(\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)\right)$. Pick $x\in\mathcal{J}(f)\setminus\left(\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)\right)$ and write $h(x)=(\alpha_{0},\alpha_{1},\dots,\alpha_{i_{0}}\dots)$. For a cylinder set
$\mathcal{C}:=\{\beta=(\beta_{0},\beta_{1},\dots):\beta_{i}=\alpha_{i}:0\leq i\leq i_{0}\}\subseteq\Sigma_{A}$
containing $h(x)$, we have that
$h^{-1}(\mathcal{C})=\bigcap_{i=0}^{i_{0}}f^{-i}(U_{i}),$
where $U_{i}\in\mathcal{U}_{\ast}$ is the disk corresponding to $\alpha_{i}$. Thus $h^{-1}(\mathcal{C})$ is open in $X$ and hence $h$ is continuous. Conversely, for an open set $\Omega\subseteq\mathcal{J}(f)\setminus\left(\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)\right)$, shrinking $\Omega$ if necessary, we can assume that $\Omega$ is contained in a nontrivial disk in $\mathcal{U}_{\ast}$. Let $\ell\geq 1$ be the smallest integer such that $f^{\circ\ell}(\Omega)$ contains at least two disks in $\mathcal{U}_{\ast}$, and let $U_{i}\in\mathcal{U}_{\ast}$ be the disk containing $f^{\circ i}(\Omega)$ for $10\leq i\leq\ell-$ 1. Considering the symbol $u_{i}$ corresponding to $U_{i}$ for each $10\leq i\leq\ell-$ 1, we conclude that
$h(\Omega)=\{\omega:=(\omega_{0},\omega_{1},\dots):\omega_{i}=u_{i}\ \text{for}\ 0\leq i\leq\ell-1\}.$
Hence $h(\Omega)$ is an open set in $\Sigma_{A}$. Thus $h^{-1}$ is continuous. ∎
3.4. Finite symbols via factoring
In this subsection, let $f$ be as in §3.3 and further assume that $f$ satisfies the homogeneous condition. Thus $f$ satisfies the assumptions in Theorem 3.4. We apply the homogeneous condition to factor the symbols in Proposition 3.6 and obtain a finite Markov result:
Proposition 3.9.
There exist a symbolic dynamical system $(\Sigma_{B},\sigma)$ of finitely many states and an injection $g:\mathcal{J}(f)\to\Sigma_{B}$ such that the following hold:
  1. (1)
    $g$ is a topological conjugation on $\mathcal{J}(f)\setminus\left(\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)\right)$ between $f$ and $\sigma$.
  2. (2)
    The complement $\Sigma_{B}\setminus g(\mathcal{J}(f))$ consists of the grand orbits of finitely many periodic cycles in $(\Sigma_{B},\sigma)$.
We will rely on a series of lemmas stated below to establish Proposition 3.9. In §3.4.1, we show that the homogeneous condition can naturally extend to the forward images of singular points. In §3.4.2, we introduce the $T$-equivalence which induces a symbolic dynamical system $(\Sigma_{B},\sigma)$ of finitely many states. We also define the factor map and show that there is a factor map $\pi$ from $\Sigma_{A}$ as in Proposition 3.6 to $\Sigma_{B}$. Then in §3.4.3, we introduce $T$-truncations of an element in $\Sigma_{B}$ to characterize the elements in $\Sigma_{B}$ and the elements in $\pi(\Sigma_{A})$. Finally, we prove Proposition 3.9 in §3.4.4.
3.4.1. Homogeneous condition
Recall Definition 3.2 for the homogeneous condition. We show that the homogeneous condition can extend to the forward orbit of $\mathrm{Sing}(f)$ in the following sense.
Lemma 3.10.
Let $\gamma\in\Gamma(f)$ and write $p:=p_{\gamma}$ for the marked point of $\gamma$. Then for any $x\in\mathrm{GO}^{*}(\gamma)$, there exist a disk $U_{x}$ containing $x$ and a scaling map $T_{x}:U_{x}\to X$ fixing $x$ such that
  1. (1)
    $f^{\circ\ell_{x}}(U_{x})\subseteq U_{p}$ and $\mathrm{diam}(f^{\circ\ell_{x}}(U_{x}))=\mathrm{diam}(U_{p})$, where $\ell_{x}\geq 0$ is the smallest integer satisfying $f^{\circ\ell_{x}}(x)=p$; and
  2. (2)
    if $x_{1},x_{2}\in\mathrm{GO}^{*}(\gamma)$ are two distinct points satisfying that $x_{2}=f^{\circ j_{0}}(x_{1})$ for some smallest integer $j_{0}\geq 1$ and $p_{\gamma}\not=f^{\circ j}(x_{1})$ for any $10\leq j\leq j_{0}-$ 1, then
    1. (a)
      $f^{\circ j_{0}}(U_{x_{1}})\subseteq U_{x_{2}}$ and $\mathrm{diam}(f^{\circ j_{0}}(U_{x_{1}}))=\mathrm{diam}(U_{x_{2}})$;
    2. (b)
      in $U_{x_{1}}$,
      $f^{\circ j_{0}}\circ T_{x_{1}}=T_{x_{2}}\circ f^{\circ j_{0}}.$
Proof.
For any $x\in\mathrm{GO}^{*}(\gamma)\cap\mathrm{Sing}(f)$, consider $\ell_{x}$, $V_{x}$ and $\varphi(\gamma)$ as in Definition 3.2 and let
$r:=\min\{\mathrm{diam}(f^{\circ\ell_{x}}(V_{x})):x\in\mathrm{GO}^{*}(\gamma)\cap\mathrm{Sing}(f)\}.$
Consider a disk $U_{p}$ containing $p$ sufficiently small so that $\mathrm{diam}(U_{p})<r$, and define $T_{p}:U_{p}\to X$ by
$T_{p}:=f^{\circ\varphi(\gamma)}.$
Then shrinking $U_{p}$ if necessary, we conclude tha $T_{p}$ is a scaling map on $U_{p}$ fixing $p$.
Now for each $x\in\mathrm{GO}^{*}(\gamma)$ with $x\not=p$, let $U_{x}\subseteq X$ be the maximal disk contained in $f^{-\ell_{x}}(U_{p})$ such that $x\in U_{x}$, and let $U^{\prime}_{x}\subseteq X$ be the maximal disk contained in $f^{-\ell_{x}}(f^{\circ\varphi(\gamma)}(U_{p}))$ such that $x\in U^{\prime}_{x}$. In what follows, we define a scaling map $T_{x}$ on $U_{x}$ so that $f\circ T_{x}=T_{f(x)}\circ f$ and $T_{x}(x)=x$, which implies the conclusion of the lemma.
Let us first consider the case that $f(x)=p$. If $x$ is a regular point, then shrinking $U_{p}$ if necessary, we have that $f$ is a scaling map on $U^{\prime}_{x}$. Then in $U_{x}$, we define
$T_{x}:=(f|_{U^{\prime}_{x}})^{-1}\circ f^{\circ(\varphi(\gamma)+1)},$
If $x$ is a singular point, observing that $U_{x}\subsetneq V_{x}$, let $T_{x}$ be given in Definition 3.2 (1). Then in this case we obtain that
$f\circ T_{x}=T_{p}\circ f.$
Moreover, observe that $T_{x}$ is a scaling map on $U_{x}$ fixing $x$.
We now deal with the case that $f(x)\not=p$, and we inductively define $T_{x}$. For $1\leq j\leq\ell_{x}$, write $x_{j}=f^{\circ j}(x)$, and suppose that there is a well-defined scaling $T_{x_{j}}:U_{x_{j}}\to X$ fixing $x_{j}$, such that for $11\leq j\leq\ell_{x}-$ 1,
  • $f\circ T_{x_{j}}=T_{x_{j+1}}\circ f$ in $U_{x_{j}}$; and
  • if $x_{j}$ is a singular point, the $T_{x_{j}}$ is given in Definition 3.2 (1).
If $x$ is a regular point, then shrinking $U_{p}$ if necessary, we have that $f$ is a scaling map on $U^{\prime}_{x}$. Then in $U_{x}$, we define
$T_{x}:=(f|_{U^{\prime}_{x}})^{-1}\circ T_{x_{1}}\circ f.$
If $x$ is a singular point, then let $T_{x}$ is given in Definition 3.2 (1). Then $T_{x}$ is a scaling map on $U_{x}$ fixing $x$. We now check $f\circ T_{x}=T_{x_{1}}\circ f$. It holds if $x$ is a regular point. In view of Definition 3.2 (2), it also holds if both $x$ and $x_{1}$ are regular points. In the remaining case that $x$ is a singular but $x_{1}$ is regular, shrinking $U_{p}$ if necessary, we have either $f^{\circ(\ell_{x}-1)}$ is a scaling map in $U^{\prime}_{x_{1}}$ or $x_{j_{0}}$ is a singular point for some smallest $12\leq j_{0}\leq\ell_{x}-$ 1. In the former case, by Definition 3.2 (1) and the induction assumption, observing that $\ell_{x_{1}}=\ell_{x}-1$, we have that in $U_{z}$,
$f^{\circ\ell_{x_{1}}}\circ f\circ T_{x}=f^{\circ\ell_{x}}\circ T_{x}=T_{p}\circ f^{\circ\ell_{x}}=T_{p}\circ f^{\circ\ell_{x_{1}}}\circ f=f^{\circ\ell_{x_{1}}}\circ T_{x_{1}}\circ f,$
which implies that $f\circ T_{x}=T_{x_{1}}\circ f$. In the latter case, shrinking $U_{p}$ if necessary, we can assume that $f^{\circ j_{0}-1}$ is a scaling map on $V^{\prime}_{x_{1}}$. By Definition 3.2 (2) and the induction assumption, we have that in $U_{x}$,
$f^{\circ(j_{0}-1)}\circ f\circ T_{x}=f^{\circ j_{0}}\circ T_{x}=T_{x_{j_{0}}}\circ f^{\circ j_{0}}=T_{x_{1}}\circ f^{\circ(j_{0}-1)}\circ f=f^{\circ(j_{0}-1)}\circ T_{x_{1}}\circ f,$
which also implies that $f\circ T_{x}=T_{x_{1}}\circ f$. ∎
For a repelling cycle $\gamma\in\Gamma(f)$ and any $x\in\mathrm{GO}^{*}(\gamma)$, recalling from §3.2 the closed disk $\overline{D}_{x}$, we can shrink the domain $U_{x}$ of $T_{x}$ so that $U_{x}\subseteq\overline{D}_{x}$. Moreover, we can further shrink $U_{x}$ so that if $U\in\mathcal{U}_{\ast}$ is not a subset of $U_{x}$ for any $x\in\cup_{\gamma\in\Gamma(f)}\mathrm{GO}^{*}(\gamma)$ and $f(U)\cap U_{x_{0}}\not=\emptyset$ for some $x_{0}\in\cup_{\gamma\in\Gamma(f)}\mathrm{GO}^{*}(\gamma)$, then $U_{x_{0}}\subseteq f(U)$. Indeed, considering the elements $V\in\mathcal{U}_{\ast}$ such that $V$ is not a subset of $\overline{D}_{x}$ for any $x\in\cup_{\gamma\in\Gamma(f)}\mathrm{GO}^{*}(\gamma)$ or $V$ is a subset of $\overline{D}_{p_{\gamma}}\setminus f^{-s_{\gamma}}(\overline{D}_{p_{\gamma}})$ for any $\gamma\in\Gamma(f)$, where $s_{\gamma}$ is the length of $\gamma$, we can choose $U_{x}$ to be the intersection of $f(V)$ for all above $V\in\mathcal{U}_{\ast}$ with $x\in f(V)$.
Thus, in the arguments that follow, we can always assume that $U_{x}$ is sufficiently small to satisfy the condition in the previous paragraph.
3.4.2. $T$-equivalence and factor map
Recall from Proposition 3.6 the collection $\mathcal{U}_{\ast}$ and the symbolic dynamical system $(\Sigma_{A},\sigma)$. We now define an equivalence relation in $\mathcal{U}_{\ast}$ induced by the maps $T_{x}:U_{x}\to X$ for all $x\in\cup_{\gamma\in\Gamma(f)}\mathrm{GO}^{*}(\gamma)$ given in Lemma 3.10.
Definition 3.11.
For two (not necessary distinct) disks $U$ and $U^{\prime}$ in $\mathcal{U}_{\ast}$, we say that $U$ and $U^{\prime}$ are $T$-equivalent if one of the following holds:
  1. (1)
    both $U$ and $U^{\prime}$ are contained in $U_{x}$ for some $x\in\cup_{\gamma\in\Gamma(f)}\mathrm{GO}^{*}(\gamma)$ and $U=(T_{x})^{\circ m}(U^{\prime})$ for some $m\in\mathbb{Z}$; or
  2. (2)
    $U=U^{\prime}$ is not contained in $U_{x}$ for any $x\in\cup_{\gamma\in\Gamma}\mathrm{GO}^{*}(\gamma)$.
Denote by $[U]$ the $T$-equivalence class for $U\in\mathcal{U}_{\ast}$ and by $\mathcal{T}$ the set formed by all the $T$-equivalence classes. The map $f$ induces a map, denoted by $\mathfrak{F}$, from $\mathcal{T}$ to its power set $2^{\mathcal{T}}$:
$\mathfrak{F}:\mathcal{T}\to 2^{\mathcal{T}}$
satisfying that for any $[U],[V]\in\mathcal{T}$, the image $\mathfrak{F}([U])$ contains $[V]$ if and only if there exist $U^{\prime},V^{\prime}\in\mathcal{U}_{\ast}$ with $[U^{\prime}]=[U]$ and $[V^{\prime}]=[V]$ such that $V^{\prime}\subseteq f(U^{\prime})$. The map $\mathfrak{F}$ admits a natural transition matrix $B$, that is, the entry of $B$ for the correspondence from $[U]$ to $[V]$ is $1$ if $[V]\in\mathfrak{F}([U])$, and is 00 if otherwise.
The $T$-equivalence classes will induce a symbolic dynamical system of finitely many states for $f$ in $\mathcal{J}(f)$ (see Lemma 3.13). Let us first take the following definition.
Definition 3.12.
For two symbolic dynamical systems $(\Sigma_{A_{1}},\sigma)$ and $(\Sigma_{A_{2}},\sigma)$ over symbolic sets $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, we say a map $\pi:\Sigma_{A_{1}}\to\Sigma_{A_{2}}$ is a factor map if the following hold:
  1. (1)
    $\pi$ is continuous,
  2. (2)
    there exists a surjection $\theta:\mathcal{A}_{1}\to\mathcal{A}_{2}$ such that $\pi((\alpha_{1},\alpha_{2},\dots))=(\theta(\alpha_{1}),\theta(\alpha_{2}),\dots)$ for any $(\alpha_{1},\alpha_{2},\dots)\in\Sigma_{A_{1}}$.
Observe that any factor map $\pi:\Sigma_{A_{1}}\to\Sigma_{A_{2}}$ satisfies that $\sigma\circ\pi=\pi\circ\sigma$ since for any $(\alpha_{1},\alpha_{2},\dots)\in\Sigma_{A_{1}}$,
$\sigma\circ\pi((\alpha_{1},\alpha_{2},\dots))=(\theta(\alpha_{2}),\dots)=\pi\circ\sigma((\alpha_{1},\alpha_{2},\dots)).$
We now introduce a factor map from $\Sigma_{A}$ to a symbolic space of finitely many states. Regard each class in $\mathcal{T}$ as a symbol and observe that the map $\theta:\mathcal{U}_{\ast}\to\mathcal{T}$, sending $U\in\mathcal{U}_{\ast}$ to $[U]\in\mathcal{T}$, is a surjection. Considering the transition matrix $B$ for $\mathfrak{F}$, we obtain a symbolic space $\Sigma_{B}$ of finitely many states. Define the map
$\pi:\Sigma_{A}\to\Sigma_{B},$
sending $(\alpha_{0},\alpha_{1},\dots)\in\Sigma_{A}$ to $(\theta(\alpha_{0}),\theta(\alpha_{1}),\dots)\in\Sigma_{B}$.
Lemma 3.13.
The map $\pi:\Sigma_{A}\to\Sigma_{B}$ is a factor map.
Proof.
From previous paragraph, we only need to show that $\pi$ is continuous. Considering any $\beta=(\beta_{0},\beta_{1},\dots)\in\Sigma_{A}$ and a sequence $\{\beta^{(m)}:=(\beta^{(m)}_{0},\beta^{(m)}_{1},\dots)\}_{m\geq 1}$ such that $\beta^{(m)}\to\beta$, as $m\to\infty$, in $\Sigma_{A}$, we have that, given $j\geq 1$, there exists $m_{j}\geq 1$ such that $\theta(\beta_{j})=\theta(\beta_{j}^{(m)})$, for all $m\geq m_{j}$. Moreover, we can choose $m_{j+1}>m_{j}$. Thus $\pi(\beta^{(m)})\to\pi(\beta)$, as $m\to\infty$. Hence $\pi$ is continuous. ∎
3.4.3. $T$-truncation
We now study the elements in $\Sigma_{B}$ and in $\pi(\Sigma_{A})$ by considering their truncations. For each $x\in\cup_{\gamma\in\Gamma}\mathrm{GO}^{*}(\gamma)$, recall that $U_{x}\subseteq\overline{D}_{x}$ is the domain of the map $T_{x}$.
Definition 3.14.
For $[\alpha]=([\alpha_{0}],[\alpha_{1}],\dots)\in\Sigma_{B}$ with $[\alpha_{i}]=[U_{i}]$ for some $U_{i}\in\mathcal{U}_{\ast}$, we say that a truncation $([\alpha_{j_{0}}],[\alpha_{j_{0}+1}],\dots,[\alpha_{j_{0}+s_{0}}])$ with $0\leq j_{0}<+\infty$ and $0\leq s_{0}\leq+\infty$ is a $T$-truncation of $[\alpha]$ if there exists a repelling cycle $\gamma\in\Gamma(f)$ such that the following hold:
  1. (1)
    for each $j_{0}\leq i\leq j_{0}+s_{0}$, $U_{i}\subset U_{x_{i}}$ for some $x_{i}\in\mathrm{GO}^{*}(\gamma)$,
  2. (2)
    if $j_{0}\geq 1$, then $U_{j_{0}-1}\cap U_{x}=\emptyset$ for any $x\in\mathrm{GO}^{*}(\gamma)$, and
  3. (3)
    if $s_{0}\not=+\infty$, then $U_{j_{0}+s_{0}+1}\cap U_{x}=\emptyset$ for any $x\in\mathrm{GO}^{*}(\gamma)$.
Moreover, if in addition $s_{0}=+\infty$, we say $([\alpha_{j_{0}}],[\alpha_{j_{0}+1}],\dots,[\alpha_{j_{0}+s_{0}}])$ is an infinite $T$-truncation; if otherwise, we say $([\alpha_{j_{0}}],[\alpha_{j_{0}+1}],\dots,[\alpha_{j_{0}+s_{0}}])$ is a finite $T$-truncation.
For a finite $T$-truncation $([\alpha_{j_{0}}],[\alpha_{j_{0}+1}],\dots,[\alpha_{j_{0}+s_{0}}])$, by the choice of $U_{x}$ for any $x\in\cup_{\gamma\in\Gamma}\mathrm{GO}^{*}(\gamma)$, each of $[\alpha_{j_{0}-1}]$ (if exists) and $[\alpha_{j_{0}+s_{0}+1}]$ is not contained in a $T$-truncation. Hence both $U_{j_{0}-1}$ (if exists) and $U_{j_{0}+s_{0}+1}$ are unique by Definition 3.11 (2). Moreover, $U_{i}$ is unique for each $j_{0}\leq i\leq j_{0}+s_{0}$:
Lemma 3.15.
Let $([\alpha_{j_{0}}],[\alpha_{j_{0}+1}],\dots,[\alpha_{j_{0}+s_{0}}])$ be a finite $T$-truncation of $([\alpha_{0}],[\alpha_{1}],\dots)\in\Sigma_{B}$. Then for each $j_{0}\leq i\leq j_{0}+s_{0}$, there exists a unique $U_{i}\in\mathcal{U}_{\ast}$ such that $[\alpha_{i}]=[U_{i}]$.
Proof.
Write $[\alpha_{i}]=[U_{i}],i\geq 0$ for some $U_{i}\in\mathcal{U}_{\ast}$, and let $\gamma$ be the repelling cycle such that $U_{i},j_{0}\leq i\leq j_{0}+s_{0},$ is a subset of $U_{x_{i}}$ for some $x_{i}\in\mathrm{GO}^{\ast}(\gamma)$. Let $p:=p_{\gamma}\in\gamma$ be marked point of $\gamma$ and consider the set $U_{p}$. It follows that $U_{j_{0}+s_{0}}\subseteq U_{p}$ and $U_{j_{0}+s_{0}+1}\cap U_{p}=\emptyset$. Since $U_{j_{0}+s_{0}+1}$ is unique and $f$ is a scaling in $U_{p}$, we conclude that $U_{j_{0}+s_{0}}$ is unique. For some $j_{0}\leq i\leq j_{0}+s_{0}$, if $U_{i}$ and $U^{\prime}_{i}$ in $\mathcal{U}_{\ast}$ satisfy $[U_{i}]=[U^{\prime}_{i}]$, since $f^{\circ j_{0}+s_{0}-i}(U_{i})=U_{j_{0}+s_{0}}=f^{\circ j_{0}+s_{0}-i}(U^{\prime}_{i})$, we conclude that $U_{i}=U^{\prime}_{i}$. Hence $U_{i}$ is unique for each $j_{0}\leq i\leq j_{0}+s_{0}$. ∎
Corollary 3.16.
Pick $[\alpha]\in\Sigma_{B}$ and assume that $[\alpha]$ has no infinite $T$-truncation. Then $[\alpha]\in\pi(\Sigma_{A})$.
Proof.
By Definition 3.11 (2) and Lemma 3.15, it suffices to check that if exactly one of $[\alpha_{j}]$ and $[\alpha_{j+1}]$ is not in a distinct $T$-truncation, then there exist $U_{j},U_{j+1}$ in $\mathcal{U}_{\ast}$ such that $[\alpha_{j}]=[U_{j}]$, $[\alpha_{j+1}]=[U_{j+1}]$ and $U_{j+1}\subseteq f(U_{j})$. If $[\alpha_{j}]$ is not in a $T$-truncation, then $U_{j}$ is unique. Moreover, by the choice of $U_{x},x\in\cup_{\gamma\in\Gamma}\mathrm{GO}^{*}(\gamma)$, for any $U_{j+1}\in\mathcal{U}_{\ast}$ with $[\alpha_{j+1}]=[U_{j+1}]$, we have $U_{j+1}\subseteq f(U_{j})$. If $[\alpha_{j=1}]$ is not in a $T$-truncation, then $U_{j+1}$ is unique, and hence there exists a desired $U_{j}$. Thus the conclusion holds. ∎
We now consider the elements in $\Sigma_{B}$ having infinite $T$-truncations. Recall from Definition 3.2 the map $\varphi$.
Lemma 3.17.
Let $[\alpha]=([\alpha_{0}],[\alpha_{1}],\dots)\in\Sigma_{B}$ and write $[\alpha_{i}]=[U_{i}]$ with $U_{i}\in\mathcal{U}_{\ast}$ for each $i\geq 0$. Then the following are equivalent:
  1. (1)
    $[\alpha]$ has an infinite $T$-truncation.
  2. (2)
    There exists a repelling cycle $\gamma:=\gamma_{[\alpha]}\in\Gamma(f)$ such that $U_{i}\subseteq\cup_{x\in\gamma}U_{x}$ for all sufficiently large $i\geq 0$.
Moreover, in this case, $[\alpha]$ eventually maps, under $\sigma$, to a periodic cycle in $\Sigma_{B}$ whose length is $\varphi(\gamma_{[\alpha]})$.
Proof.
If there exists such a repelling cycle $\gamma$, let $m_{0}$ be the smallest integer such that $U_{m_{0}}$ is a subset of $\cup_{x\in\gamma}U_{x}.$ Given any large $i\geq 0$, we can choose $U^{\prime}_{m_{0}}\subseteq\cup_{x\in\gamma}U_{x}$ such that $[U^{\prime}_{m_{0}}]=[U_{m_{0}}]$ and $f^{i-m_{0}}(U^{\prime}_{m_{0}})$ is also a subset of $\cup_{x\in\gamma}U_{x}$. Note that $[f^{i-m_{0}}(U^{\prime}_{m_{0}})]=[U_{i}]$. Thus $[\alpha]$ has a $T$-truncation containing $[\alpha_{m_{0}}],[\alpha_{m_{0}+1}],\dots,[\alpha_{i}]$. Since $i$ is arbitrarily large, we conclude that $[\alpha]$ has an infinite $T$-truncation.
If there is no such a repelling cycle $\gamma$, then there exists a subsequence $\{[\alpha_{i_{j}}]\}_{j=1}^{\infty}$ such that exactly one of the following holds:
  1. (1)
    each $[\alpha_{i_{j}}]$ is not contained in a $T$-truncation of $[\alpha]$, or
  2. (2)
    for each $j\geq 1$, there exist two distinct repelling cycles $\gamma_{j}$ and $\gamma_{j+1}$ in $\Gamma(f)$ with $U_{i_{j}}\subseteq\cup_{x\in\gamma_{i}}U_{x}$ and $U_{i_{j+1}}\subseteq\cup_{y\in\gamma_{i+1}}U_{y}$.
In either case, $[\alpha]$ has no infinite $T$-truncations.
Moreover, if $[\alpha]$ has an infinite $T$-truncation, then $[\alpha_{i}]=[\alpha_{i+\varphi(\gamma)}]$ for sufficiently large $i\geq 0$, since $T_{x}=f^{\circ\varphi(\gamma_{[\alpha]})}$ for any $x\in\gamma_{[\alpha]}$ (see Lemma 3.10). Thus $[\alpha]$ eventually maps, under $\sigma$, to a cycle in $\Sigma_{B}$ of length $\varphi(\gamma_{[\alpha]})$. ∎
Corollary 3.18.
There are only finitely many periodic elements in $\Sigma_{B}$ having infinite $T$-truncations.
Proof.
Since $\mathcal{T}$ is finite, given an integer $s\geq 1$, there are only finitely many periodic cycles in $\Sigma_{B}$ of length $s$. Thus the conclusion follows from Lemma 3.17. ∎
We now identify the points in $\Sigma_{A}$ whose images, under the factor map $\pi:\Sigma_{A}\to\Sigma_{B}$, contain infinite $T$-truncations to the points in $\mathcal{J}(f)$ that are eventually mapped to repelling cycles in $\Gamma(f)$. Recall from Proposition 3.6 the map $h:\mathcal{J}(f)\to\Sigma_{A}$.
Lemma 3.19.
Let $\alpha\in\Sigma_{A}$ such that $\pi(\alpha)$ has an infinite T-truncation. Then
$h^{-1}(\alpha)\in\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma).$
Proof.
Suppose on the contrary that $h^{-1}(\alpha)$ is not contained in $\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)$. Write $\pi(\alpha)=([\alpha_{0}],[\alpha_{1}],\dots)$, and let $U_{i}\in\mathcal{U}_{\ast}$ such that $[\alpha_{i}]=[U_{i}]$ for each $i\geq 0$. Consider any $T$-truncation $([\alpha_{j_{0}}],[\alpha_{j_{0}+1}],\dots,[\alpha_{j_{0}+s_{0}}])$ of $[\alpha]$. Let $\gamma\in\Gamma(f)$ with marked point $p:=p_{\gamma}$ such that $U_{i}\subseteq U_{x_{i}}$ with $x_{i}\in\mathrm{GO}^{*}(\gamma)$ for each $j_{0}\leq i\leq j_{0}+s_{0}$. Then there exists $j_{0}\leq m_{0}\leq j_{0}+s_{0}$ such that $U_{m_{0}}\subseteq U_{p}$ but $f(U_{m_{0}})\cap U_{x}=\emptyset$ for any $x\in\mathrm{GO}^{*}(\gamma)$, since $U_{p}\subseteq\overline{D}_{p}$, where $\overline{D}_{p}$ is as in §3.2. Thus $m_{0}=j_{0}+s_{0}$ and hence $s_{0}\not=+\infty$. It follows that $([\alpha_{j_{0}}],[\alpha_{j_{0}+1}],\dots,[\alpha_{j_{0}+s_{0}}])$ is a finite $T$-truncation. This contradicts that $\pi(\alpha)$ has an infinite $T$-truncation. Thus the conclusion follows. ∎
3.4.4. Proof of Proposition 3.9
To establish Proposition 3.9, we first show the injectivity of the factor map $\pi:\Sigma_{A}\to\Sigma_{B}$.
Lemma 3.20.
The factor map $\pi:\Sigma_{A}\to\Sigma_{B}$ is injective.
Proof.
Pick $[\alpha]=([\alpha_{0}],[\alpha_{1}],\dots)\in\pi(\Sigma_{A})$. If $[\alpha]$ has an infinite $T$-truncation, then by Lemma 3.19, there is a unique $\alpha\in\Sigma_{A}$ corresponding to a point in $\cup_{\gamma\in\Gamma(f)}\mathrm{GO}(\gamma)$ such that $\pi(\alpha)=[\alpha]$. We now assume that $[\alpha]$ has no infinite $T$-truncation. If $[\alpha_{i}]$ is not in a $T$-truncation of $[\alpha]$, then by Definition 3.11 (2), there has a unique $U_{i}\in\mathcal{U}_{i}$ such that $[\alpha_{i}]=[U_{i}]$. If $[\alpha_{i}]$ is in a (finite) $T$-truncation of $[\alpha]$, then by Lemma 3.15, there also has a unique $U_{i}\in\mathcal{U}_{i}$ such that $[\alpha_{i}]=[U_{i}]$. Thus $\pi^{-1}([\alpha_{i}])$ is a singleton. ∎
Now let us prove Proposition 3.9.
Proof of Proposition 3.9.
Let $h$ be as in Proposition 3.6 and consider the factor map $\pi:\Sigma_{A}\to\Sigma_{B}$ as in Lemma 3.13. Define $g=\pi\circ h$. By Proposition 3.6 and Lemma 3.20, we have that $g$ is injective.
Statement (1) follows immediately from Proposition 3.6, Lemmas 3.13 and 3.20. For statement (2), it follows from Corollary 3.16 that each element $\Sigma_{B}\setminus g(\mathcal{J}(f))$ has an infinite $T$-truncation. Then Lemmas 3.17 and 3.19 imply that $\Sigma_{B}\setminus g(\mathcal{J}(f))$ is non-empty and each element is eventually mapped to a periodic cycle. Moreover, there are finitely many such periodic cycles by Corollary 3.18. Thus statement (2) holds. ∎
3.5. Proof of Theorem 3.4
Proposition 3.9 implies that for each $n\geq 1$, the set $\mathrm{Fix}^{*}(f^{\circ n})$ is finite, since $\sigma^{\circ n}$ has finitely many fixed points in $\Sigma_{B}$. Moreover, it follows from Proposition 3.9 (1) and Lemmas 3.17 and 3.19 that the image $g(\gamma)$ of a repelling cycle $\gamma$ in $\mathcal{J}(f)$ of length $s\geq 1$ is a periodic cycle of $\sigma$ in $\Sigma_{B}$ of length $s$.
Let $\beta_{1},\dots,\beta_{m_{0}}$ be all the periodic cycles in $\Sigma_{B}\setminus g(\mathcal{J}(f))$, and let $\ell_{m}$ be the length of $\beta_{m}$ for each $1\leq m\leq m_{0}$. It follows from Corollaries 3.16 and 3.18 that $m_{0}$ is finite. We conclude that
$\zeta_{f}(X,t)=\frac{\zeta_{\sigma}(\Sigma_{B},t)}{\exp\left(\sum_{m=1}^{m_{0}}\left(\sum_{j=1}^{\infty}\left(\ell_{m}\frac{t^{(j\ell_{m})}}{j\ell_{m}}\right)\right)\right)},$
where $\zeta_{\sigma}(\Sigma_{B},t)$ is the Artin-Mazur zeta function for $(\Sigma_{B},\sigma)$. In view of §2.2,
$\zeta_{\sigma}(\Sigma_{B},t)=\det(I-tB)^{-1}.$
Noting that
$\exp\left(\sum_{m=1}^{m_{0}}\left(\sum_{j=1}^{\infty}\left(\ell_{m}\frac{t^{(j\ell_{m})}}{j\ell_{m}}\right)\right)\right)=\prod_{m=1}^{m_{0}}\frac{1}{1-t^{\ell_{m}}},$
we conclude that
$\zeta_{f}(X,t)=\frac{\prod_{m=1}^{m_{0}}(1-t^{\ell_{m}})}{\det(I-tB)}.$
Let $G(t):=\gcd(\prod_{m=1}^{m_{0}}(1-t^{\ell_{m}}),\det(I-tB))$. If $G(t)\not=1$, then all the zeros of $G(t)$ are roots of unity. It follows that both $(\prod_{m=1}^{m_{0}}(1-t^{\ell_{m}}))/G(t)$ and $\det(I-tB)/G(t)$ are in $\mathbb{Z}[t]$. Thus $\zeta_{f}(X,t)\in\mathbb{Z}(t)$. ∎
4. Rational maps case
The goal of this section is to prove Theorems 1.1 and 1.2. We establish Theorem 1.1 in §4.1 and Theorem 1.2 in §4.2, respectively.
4.1. Proof of Theorem 1.1
In view of §2.1 and Theorem 3.4, to establish Theorem 1.1, it suffices to show that $\phi$ satisfies the normality of critical points and the homogeneous condition. The normality of critical points follows immediately from [3, Theorem 3.5] and [9, Theorem D]. Now let us show that $\phi$ satisfies the homogeneous condition. We can regard $\phi\in K(z)$ as a rational map in $K^{a}(z)$, where $K^{a}$ is an algebraic closure of $K$, and denote by $m_{x}(\phi)$ the local degree of $\phi\in K^{a}(z)$ for any $x\in\mathbb{P}^{1}(K^{a})$. Recall the set $\Gamma(\phi)$ of repelling cycles in $\mathcal{J}(\phi,K)$, each of which has grand orbit containing a critical point in $\mathrm{Crit}(\phi,K)$.
Lemma 4.1.
Any subhyperbolic rational map $\phi\in K(z)$ of degree at least $2$ satisfies the homogeneous condition.
Proof.
For any $\gamma\in\Gamma(\phi)$ with a marked point $p_{\gamma}$, set
$\delta_{\gamma}:=s_{\gamma}\cdot\prod_{c\in\mathrm{GO}(\gamma)\cap\mathrm{Crit}(\phi,K)}m_{c}(\phi),$
where $s_{\gamma}$ is the length of $\gamma$. We define the map
$\begin{array}[]{cccc}\varphi:&\Gamma(\phi)&\to&\mathbb{Z}_{>0}\\ &\gamma&\to&\delta_{\gamma}.\end{array}$
Pick $c\in\mathrm{GO}(\gamma)\cap\mathrm{Crit}(\phi,K)$ and let $\ell_{c}\geq 1$ be the smallest integer such that $\phi^{\circ\ell_{c}}(c)=p_{\gamma}$. Conjugating $\phi$ by a Möbius transformation in $K(z)$, we can assume that $c\in K$ and $p=0$. Then in a small neighborhood $U_{p}\subseteq K$ of $p$, we can write
$\phi^{\circ\delta_{\gamma}}(z)=\lambda^{\delta_{\gamma}/s_{\gamma}}z+\sum_{i=2}^{\infty}a_{i}z^{i},$
where $\lambda=(\phi^{\circ s_{\gamma}})^{\prime}(0)$ and $a_{i}\in K$ for each $i\geq 2$. Moreover, in a sufficiently small neighborhood $U_{c}\subseteq K$ of $c$ with $\phi^{\circ\ell_{c}}(U_{c})\subseteq U_{p}$, we can write
$\phi^{\circ\ell_{c}}(z)=\sum_{i=0}^{\infty}b_{i}(z-c)^{m+i},$
where $b_{0}\in K\setminus\{0\}$, $b_{i}\in K$ for each $i\geq 1$, and $m=m_{c}(\phi^{\circ\ell_{c}})$.
Shrinking $U_{c}$ if necessary, we claim that, in $U_{c}$, there exists
$T_{c}(z)=c+\sum_{i=1}^{\infty}\alpha_{i}(z-c)^{i}$
with $\alpha_{1}\in K\setminus\{0\}$ and $\alpha_{i}\in K$ for each $i\geq 2$ such that
$\phi^{\circ\ell_{c}}\circ T_{c}=\phi^{\circ(\delta_{\gamma}+\ell_{c})}.$
Observing that $m$ divides $\delta_{\gamma}/s_{\gamma}$, we can take
$\alpha_{1}=\lambda^{\delta_{\gamma}/(s_{\gamma}\cdot m)}\in K\setminus\{0\},$
and inductively, for each $i\geq 1$, comparing the coefficients of $(z-c)^{m+i}$ in $\phi^{\circ\ell_{c}}\circ T_{c}$ and in $\phi^{\delta_{\gamma}+\ell_{c}}$, we obtain that
$b_{0}\alpha_{1}^{m-1}\alpha_{i+1}=P(a_{2},\dots,a_{i},b_{0},\dots,b_{i},\alpha_{1},\dots,\alpha_{i}),$
where $P\in\mathbb{Z}[z_{1},\dots,z_{3i}]$ is a polynomial, and conclude that $\alpha_{i+1}\in K$ is unique. Thus $T_{c}$ is unique under the choice of $\alpha_{1}$. Further shrinking $U_{c}$ if necessary, we conclude that $T_{c}$ is a scaling in $U_{c}$.
To complete the proof, we need to consider the case that $\mathrm{GO}(\gamma)\cap\mathrm{Crit}(\phi,K)$ contains two distinct critical points, say $c_{1}$ and $c_{2}$, with $\phi^{\circ j_{0}}(c_{1})=c_{2}$ for some smallest $j_{0}\geq 1$. We must check $\phi^{\circ j_{0}}\circ T_{c_{1}}=T_{c_{2}}\circ\phi^{\circ j_{0}}$. By previous paragraph, we only need to show that there exists a scaling map $f$, defined near $c_{1}$,fixing $c_{1}$ and having $f^{\prime}(c_{1}):=\lambda^{\delta_{\gamma}/(s_{\gamma}\cdot m_{1})}z$, where $m_{1}=m_{c_{1}}(\phi^{\circ\ell_{c_{1}}})$, such that $\phi^{\circ j_{0}}\circ f=T_{c_{2}}\circ\phi^{\circ j_{0}}$; indeed, such a map $f$ satisfies $\phi^{\circ\ell_{c_{1}}}\circ f=\phi^{\circ(\delta_{\gamma}+\ell_{c_{1}})}$, which implies that $f=T_{c_{1}}$ by the uniqueness of $T_{c_{1}}$. Post-and pre-composing translations in $K[z]$, we can assume that $c_{1}=c_{2}=0$ and that
$\phi^{j_{0}}(z)=\sum_{j\geq 0}B_{j}z^{d+j},$
where $d=m_{c_{1}}(\phi^{j_{0}})$, $B_{0}\in K\setminus\{0\}$ and $B_{j}\in K$ for $j\geq 1$. From the previous paragraph, we have that
$T_{c_{2}}(z)=\lambda^{\delta_{\gamma}/(s_{\gamma}\cdot m_{2})}z+\sum_{j\geq 2}\beta_{j}z^{j},$
where $m_{2}=m_{c_{2}}(\phi^{\circ\ell_{c_{2}}})$ and $\beta_{j}\in K$ for $j\geq 2$. We claim there exists a desired $f(z)=\sum_{j\geq 1}A_{j}z^{j}$. To determine $A_{j}$, we compare the coefficients of $\phi^{\circ j_{0}}\circ f$ and $T_{c_{2}}\circ\phi^{\circ j_{0}}$. First, from the coefficients of $z^{d}$, we obtain
$A_{1}^{d}=\lambda^{\delta_{\gamma}/(s_{\gamma}\cdot m_{2})}.$
Since $d\cdot m_{2}=m_{1}$ and $m_{1}$ divides $\delta_{\gamma}/s_{\gamma}$, we can let
$A_{1}=\lambda^{\delta_{\gamma}/(s_{\gamma}\cdot m_{1})}.$
For each $j\geq 1$, comparing the coefficients of $z^{d+j}$, we obtain that
$B_{0}A_{1}^{d-1}A_{j+1}=Q(A_{1},\dots,A_{j},B_{0},\dots,B_{j},\beta_{2},\dots,\beta_{j}),$
where $Q\in\mathbb{Z}[z_{1},\dots,z_{3j}]$ is a polynomial, which implies that $A_{j+1}\in K$ exists. This shows the existence of $f$ and finishes the proof. ∎
Lemma 4.1 completes the proof of Theorem 1.1. ∎
4.2. Entropy
We aim to establish Theorem 1.2 in this section. From previous sections, the main part is to construct a subhyperbolic rational map in $K(z)$ whose entropy is the logarithm of a given weak Perron number $\lambda$. As aforementioned in §1, we first construct a matrix with leading eigenvalue $\lambda$ satisfying the properties that reflect the non-archimedean property, see §4.2.2. The construction replies on some preliminaries in the graph theory introduced in §4.2.1. Moreover, the resulting matrix induces a piecewise linear map and records the transition of the pieces in the domain, see §4.2.3. Then, in §4.2.4, we glue the pieces to obtain a hyperbolic rational map in $K(z)$ and study its dynamics. In §4.2.5, we show Theorem 1.2 and provide an example to illustrate our construction. In this section, we take the field $K$ as in Theorem 1.2 and let $\mathfrak{u}\in K$ be a uniformizer of $K$. To simplify the notation, when $K$ is clear from the context, we write ||$\cdot$|| ⋅ | instead of $|\cdot|_{K}$. Moreover, a closed disk $\overline{D}$ in $K$ is rational if $\mathrm{diam}(\overline{D})$ is contained in the value group $|K\setminus\{0\}|$.
4.2.1. Augmented graphs
Let $G=(V,E)$ be a finite directed graph, where $V:=V(G)$ is the set of vertices of $G$ and $E:=E(G)$ is the set of directed edges of $G$. Recall from §2.2 the transition matrix $A_{G}$ of $G$. We say that $G$ is simple if there is at most one directed edge with initial point $v$ and end point $v^{\prime}$ for any two given vertices $v,v^{\prime}\in V$; equivalently, $A_{G}$ is a (0,1)-matrix. One way to obtain a simple graph from a given finite directed graph $G_{0}$ is to construct augmented graphs by adding extra vertices:
Definition 4.2.
Given an integer $n\geq 1$, an $n$-augmented graph $G$ of $G_{0}$ is a finite directed graph obtained by replacing each $e\in E(G_{0})$ with a directed path of length $n$ via adding $n-1$ extra vertices $v^{e}_{1},\dots,v^{e}_{n-1}$ in between with $v^{e}_{i}\not=v^{e^{\prime}}_{j}$ for any directed edges $e\not=e^{\prime}$ of $G_{0}$ and any $11\leq i\leq j\leq n-$ 1.
If $G$ is an $n$-augmented graph of $G_{0}$, then #$\#E(G)=n\cdot\#E(G_{0})$ and $V(G_{0})$ is naturally identified to a subset of $V(G)$. Moreover, for each vertex $v^{e}_{i},1\leq i\leq n$ of $G$, there exist a unique $e_{1}\in E(G)$ having initial point $v^{e}_{i}$ and a unique $e_{2}\in E(G)$ having end point $v^{e}_{i}$, where $v^{e}_{i}$ is as in Definition 4.2.
The following result states the leading eigenvalue of the transition matrix of an augmented graph. Recall from §2.2 the Artin-Mazur zeta function $\zeta(G,t)=\det(I-tA_{G})^{-1}$.
Lemma 4.3.
Let $G_{0}$ be a finite directed graph with $\lambda_{0}\geq 1$ as the leading eigenvalue of $A_{G_{0}}$, and given $n\geq 1$, let $G$ be an $n$-augmented graph of $G_{0}$. Then the leading eigenvalue $\lambda\geq 1$ of $A_{G}$ satisfies that $\lambda^{n}=\lambda_{0}$.
Proof.
Observe that for any $m\geq 1$, the number $N(G_{0},m)$ of loops of length $m$ in $G_{0}$ equals the number of loops of length $nm$ in $G$. Thus
$\det(I-tA_{G})^{-1}=\zeta(G,t)$ $=\exp\left(\sum_{m=1}^{\infty}N(G_{0},m)\frac{t^{mn}}{nm}\right)=\exp\left(\frac{1}{n}\sum_{m=1}^{\infty}N(G_{0},m)\frac{(t^{n})^{m}}{m}\right)$
$=\exp\left(\frac{1}{n}\log\left(\det(I-t^{n}A_{G_{0}})^{-1}\right)\right)=\det(I-t^{n}A_{G_{0}})^{-1/n}.$
It follows that the leading eigenvalue $\lambda\geq 1$ of $A_{G}$ is an $n$-th root of $\lambda_{0}$. ∎
4.2.2. Admissible matrices
Consider an $n\times n$ matrix $A=(a_{ij})$ and for each $1\leq i\leq n$, set
$\mathcal{I}_{A}(i):=\{1\leq j\leq n,a_{ij}\not=0\}.$
Definition 4.4.
An $n\times n$ irreducible matrix $A$ is admissible if it satisfies the following properties:
  1. (1)
    (non-zero property) there are no rows or columns that are identically zero;
  2. (2)
    (constant property) non-zero entries in each row are constant;
  3. (3)
    (containing property) if $\mathcal{I}_{A}(i_{1})\cap\mathcal{I}_{A}(i_{2})\not=\emptyset$ for some $1\leq i_{1}<i_{2}\leq n$, then either $\mathcal{I}_{A}(i_{1})\subseteq\mathcal{I}_{A}(i_{2})$ or $\mathcal{I}_{A}(i_{2})\subseteq\mathcal{I}_{A}(i_{1})$; and
  4. (4)
    (Markov property) for each $1\leq i\leq n$, there exists an integer $m_{i}\geq 1$ such that the $i$-th row of $A^{m_{i}}$ contains at least two non-zero entries.
Every weak Perron number can be archived as the leading eigenvalue of an admissible $(0,1)$-matrix:
Proposition 4.5.
Let $\lambda>1$ be a weak Perron number. Then there exists an admissible $(0,1)$-matrix with leading eigenvalue $\lambda$.
Proof.
Let $n_{0}\geq 1$ be an integer such that $\lambda^{n_{0}}$ is a Perron number. It follows from Lind’s theorem [21, Theorem 1] that there exists a non-negative integer matrix $B$ such that
  • the leading eigenvalue of $B$ is $\lambda^{n_{0}}$, and
  • all entries of $B^{j_{0}}$ are positive integers for some $j_{0}\geq 1$.
Observe that $B^{j_{0}}$ has leading eigenvalue $\lambda^{n_{0}j_{0}}$.
Let $G_{0}$ be the transition graph of $B^{j_{0}}$. Consider an $(n_{0}j_{0})$-augmented graph $G$ of $G_{0}$ and the transition matrix $A_{G}$. By Lemma 4.3, we conclude that $A_{G}$ has leading eigenvalue $\lambda$. Observing that $A_{G}$ is a $(0,1)$-matrix, to complete the proof, we show that $A_{G}$ is admissible.
Noting that $G_{0}$ is strongly connected, we conclude that any vertex of $G$ is an initial point and an end point of (possibly distinct) edges of $G$. Thus $A_{G}$ satisfies the non-zero property. Since the non-zero entries of $A_{G}$ are $1$, the constant property of $A_{G}$ holds. Let us check the containing property for $A_{G}$. For any vertex $v\in V(G)$, let $e_{1},e_{2}\in E(G)$ such that the end points of $e_{1}$ and of $e_{2}$ are $v$. Denote by $v_{1}$ the initial point of $e_{1}$ and by $v_{2}$ the initial point of $e_{2}$. If $v_{1}\not=v_{2}$, then $v\in V(G_{0})\subseteq V(G)$. It follows that $e_{1}$ (resp. $e_{2}$) is the only edge with initial point $v_{1}$ (resp. $v_{2}$). Thus $A_{G}$ satisfies the containing property. Finally, since $A_{G_{0}}=B^{j_{0}}$ satisfies the Markov property, we conclude that $A_{G}$ also satisfies the Markov property. ∎
4.2.3. Piecewise linear maps
Given an admissible $n\times n$ (0,1)-matrix $A=(a_{ij})$, we now construct a piecewise linear map whose transition matrix is $A$ in the sense of Definition 4.7 below.
Observe that $\mathcal{I}_{A}(i)$ induces equivalence relations on $L_{A}:=\{1,2,\dots,n\}$ as follows:
  • Two elements in $L_{A}$ are equivalent of level $1$ if they are contained in $\mathcal{I}_{A}(i)$ for some $i\in L_{A}$. Denote by $[j]_{1}$ the equivalence class of level $1$ containing $j\in L_{A}$. Moreover, we call $i\in L_{A}$ an indicator of $[j]_{1}$ if $j\in\mathcal{I}_{A}(i)$ and #$\#\mathcal{I}_{A}(i)=\max\{\#\mathcal{I}_{A}(i^{\prime}):j\in\mathcal{I}_{A}(i^{\prime})\}$.
  • Inductively, two elements $j_{1},j_{2}\in L_{A}$ are equivalent of level $\ell\geq 2$ if $[j_{1}]_{\ell-1}=[j_{2}]_{\ell-1}$ and $j_{1},j_{2}\in\mathcal{I}_{A}(i)$ for some $i\in L_{A}$ not an indicator of any $[j_{1}]_{s}=[j_{2}]_{s}$ for $11\leq s\leq\ell-$ 1. In this case, we say that $[j_{1}]_{\ell}$ is a successor of $[j_{1}]_{\ell-1}$. Moreover, we call $i\in L_{A}$ an indicator of $[j]_{\ell}$ if $j\in\mathcal{I}_{A}(i)$ and #$\#\mathcal{I}_{A}(i)=\max\{\#\mathcal{I}_{A}(i^{\prime}):j\in\mathcal{I}_{A}(i^{\prime})\ \text{and}\ i^{\prime}\ \text{is not an indicator of}\ [j]_{s}\ \text{for any}\ 1\leq s\leq\ell-1\}$.
Denote by $\mathcal{E}_{A}$ the set of all equivalence classes induced by the above equivalent relations. If $[j]_{\ell}\in\mathcal{E}_{A}$ but $[j]_{\ell+1}$ does not exist, we call $j$ a terminal of level $\ell$. Denote by $\mathcal{T}_{A}([j]_{\ell})$ the set of terminals $j^{\prime}$ of level $\ell$ such that $[j^{\prime}]_{\ell}=[j]_{\ell}$. For $[j]_{\ell}\in\mathcal{E}_{A}$, denote by $\mathcal{S}_{A}([j]_{\ell})$ the set of successors of $[j]_{\ell}$. A class in $\mathcal{E}_{A}$ is minimal if it has no successors. Observe that if $[j]_{\ell}\in\mathcal{E}_{A}$ is minimal, then $j$ is a terminal of level $\ell$. Moreover, each $j\in L_{A}$ is a terminal of some level.
Example 4.6.
Consider the admissible (0,1)-matrix $A=\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&0\end{bmatrix}.$ Then $[1]_{1}=[2]_{1}=[3]_{1}$ has indicators $1$ and $2$, and $[1]_{2}=[2]_{2}$ is minimal and has indicator $3$. So $\mathcal{E}_{A}=\{[1]_{1},[1]_{2}\}$ and $\mathcal{S}_{A}([1]_{1})=\{[1]_{2}\}$. Moreover, $3\in\mathcal{T}_{A}([1]_{1})$ is a terminal of level $1$, and both $1$ and $2$ are terminals in $\mathcal{T}_{A}([1]_{2})$ of level $2$.
Although the indicators for a given class in $\mathcal{E}_{A}$ may not be unique, they induce a well-defined map
$\kappa_{A}:L_{A}\to\mathcal{E}_{A},$
sending $i\in L_{A}$ to the class in $\mathcal{E}_{A}$ that has an indicator $i$. We now use $\kappa_{A}$ to construct a natural piecewise linear map in the following sense:
Definition 4.7.
Given an $n\times n$ (0,1)-matrix $A=(a_{ij})$, a map $f:\cup_{i=1}^{n}\overline{D}_{i}\to K$ defined on $n$ distinct rational closed disks $\overline{D}_{i},1\leq i\leq n$ in $K$ is linearly compatible to $A$ if the following hold:
  1. (1)
    $f|_{\overline{D}_{i}}$ is linear for each $1\leq i\leq n$, and
  2. (2)
    up to reindex $1\leq i\leq n$, $\overline{D}_{j}\subseteq f(\overline{D}_{i})$ if and only if $a_{ij}=1$ for $1\leq i,j\leq n$.
Lemma 4.8.
Let $A$ be an admissible $n\times n$ (0,1)-matrix. Then there exists a map $f:\cup_{i=1}^{n}\overline{D}_{i}\to K$ linearly compatible to $A$. Moreover, we can choose $f$ so that the following hold:
  1. (1)
    $\overline{D}_{i}\subsetneq\overline{D}(0,1)$ and $\overline{D}_{i}\cap\overline{D}(0,|\mathfrak{u}|)=\emptyset$ for each $1\leq i\leq n$.
  2. (2)
    $\cup_{j\in\mathcal{I}_{A}(i)}\overline{D}_{j}\subsetneq f(\overline{D}_{i})$ for each $1\leq i\leq n$.
Proof.
Let $N\geq 1$ be the number of classes of level $1$ in $\mathcal{E}:=\mathcal{E}_{A}$. We consider $N$ pairwise disjoint rational closed disks $U_{i},1\leq i\leq N$ in $K$, each corresponding to an equivalence class in $\mathcal{E}$ of level $1$. Since the residue field of $K$ has at least two elements, we can choose $U_{i}\subsetneq\overline{D}(0,1)$ so that $U_{i}\cap\overline{D}(0,|\mathfrak{u}|)=\emptyset$ for each $1\leq i\leq N$.
Inductively, in the rational closed disk $U$ corresponding to an element $[j]_{\ell}\in\mathcal{E}$, we consider #$\#\mathcal{S}_{A}([j]_{\ell})+\#\mathcal{T}_{A}([j]_{\ell})$ many pairwise disjoint rational closed subdisks so that
  • #$\#\mathcal{S}_{A}([j]_{\ell})$ disks one-to-one correspond to the successors in $\mathcal{S}_{A}([j]_{\ell})$;
  • the remaining #$\#\mathcal{T}_{A}([j]_{\ell})$ disks one-to-one correspond to the terminals in $\mathcal{T}_{A}([j]_{\ell})$; and
  • the union of disks corresponding to the terminals is properly contained in $U$.
Thus, we totally obtain #$\#\mathcal{E}+n$ many pairwise disjoint rational closed disks with $\#\mathcal{E}$ disks one-to-one corresponding to the classes in $\mathcal{E}$ and $n$ disks one-to-one corresponding to the terminals. Denote by $\mathcal{U}$ the set of $\#\mathcal{E}$ disks for $\mathcal{E}$, and let $\mathcal{A}$ be the set formed by the $n$ disks for the terminals.
Moreover, we get natural bijections
$\iota:\mathcal{E}\to\mathcal{U},$
sending each class in $\mathcal{E}$ to its corresponding rational closed disk, and
$\mathcal{L}:L_{A}\to\mathcal{A},$
sending $j\in\mathcal{L}$ to the rational closed disk for $j$ as a terminal.
Now recall the map $\kappa:=\kappa_{A}:L_{A}\to\mathcal{E}$. For each $i\in L_{A}$, considering the rational closed disks $\mathcal{L}(i)\in\mathcal{A}$ and $\kappa(i)\in\mathcal{E}$, we let
$f_{i}:\mathcal{L}(i)\to\iota\circ\kappa(i)$
be any linear and bijective map. Write $\overline{D}_{i}:=\mathcal{L}(i)$ and consider
$f:\cup_{i=1}^{n}\overline{D}_{i}\to K$
such that $f|_{\overline{D}_{i}}=f_{i}$ for each $1\leq i\leq n$. Then $f$ is a map linearly compatible to $A$ satisfying the desired properties. ∎
4.2.4. Gluing
Let $f:\cup_{i=1}^{n}\overline{D}_{i}\to K$ be a map in Lemma 4.8 for an admissible $n\times n$ (0,1)-matrix $A$. We now glue the pieces of $f$ to obtain hyperbolic rational map in $K(z)$ whose Julia set is contained in the domain of $f$. The gluing process employs the following reformulation of [24, Theorem 4.2]. Denote by $\mathbb{C}_{K}$ be the completion of an algebraical closure of $K$.
Lemma 4.9.
Let $B_{1},B_{2},\dots,B_{s}$ be pairwise disjoint rational closed disks in $\mathbb{C}_{K}$. For each $1\leq i\leq s$, pick $a_{i}\in B_{i}$ and set $\rho_{i}:=\min_{i\not=j}\{|a_{i}-a_{j}|\}$. Choose $b_{i}\in\mathbb{C}_{K}$ such that $\mathrm{diam}(B_{i})<|b_{i}|<\rho_{i}$. Consider rational maps $g_{1},g_{2},\dots,g_{s}$ in $\mathbb{C}_{K}(z)$ such that
  1. $\mathrm{(i)}$
    $g_{i}(B_{i})$ is a disk in $\mathbb{C}_{K}$ for each $1\leq i\leq s$; and
  2. $(\mathrm{ii})$
    $g_{i}(\cup_{i=1}^{s}B_{i})$ is uniformly bounded for all $1\leq i\leq s$.
Then for any $\epsilon>0$, there exists $M_{0}>1$ such that for all $M>M_{0}$, the map
$F(z)=\sum_{i=1}^{s}g_{i}(z)h_{i}(z),$
where
$h_{i}(z)=\frac{1}{1-\left(\frac{z-a_{i}}{b_{i}}\right)^{M}}$
satisfies the following: for each $1\leq i\leq s$,
  1. (1)
    $F(B_{i})=g_{i}(B_{i})$ and in fact $|F(z)-g_{i}(z)|<\epsilon$ for any $z\in B_{i}$; and
  2. (2)
    $\deg_{B_{i}}F=\deg_{B_{i}}g_{i}$
Set $\overline{D}_{0}=\overline{D}(0,|\mathfrak{u}|)$ and define
$f_{0}(z):\overline{D}_{0}\to\overline{D}_{0}$
to be the identity map. For $1\leq i\leq n$, write
$f_{i}:=f|_{\overline{D}_{i}}.$
Note that the maps $f_{i},0\leq i\leq n,$ are in fact well-defined on $\mathbb{C}_{K}$. For each $0\leq i\leq n$, pick $a_{i}\in\overline{D}_{i}$ and choose $x_{i}\in K$ such that $|x_{i}|=\mathrm{diam}(\overline{D}_{i})$. Set
$b_{i}:=\frac{x_{i}}{\sqrt{\mathfrak{u}}}.$
Then each $b_{i}$ is contained in the quadratic extension $K(\sqrt{\mathfrak{u}})$ of $K$. Let $M\geq 1$ be a sufficiently large even number to be determined later. For each $0\leq i\leq n$, we consider
$h_{i}(z):=\frac{1}{1-\left(\frac{z-a_{i}}{b_{i}}\right)^{M}}.$
and define
$F(z):=\sum_{i=0}^{n}f_{i}(z)h_{i}(z).$
Since $M$ is even, we conclude that $h_{i}\in K(z)$ and hence $F\in K(z)$.
Proposition 4.10.
For a sufficiently large even integer $M\geq 1$, the map $F$ is a hyperbolic map such that the following hold:
  1. (1)
    $\mathcal{J}(F,K)\not=\emptyset$ is compact.
  2. (2)
    $\mathcal{J}(F,K)\subseteq\cup_{i=1}^{n}\overline{D}_{i}$ and $\mathcal{J}(F,K)\cap\overline{D}_{i}\not=\emptyset$ for each $1\leq i\leq n$.
  3. (3)
    The transition matrix of $F$ on $\overline{D}_{i},1\leq i\leq n,$ is $A$
The proof of Proposition 4.10 is provided after we state the following two lemmas on the Fatou set and Julia set of $F$.
Lemma 4.11.
There exists a sufficiently large even integer $M\geq 1$ such that
$F(\mathbb{P}^{1}(K)\setminus(\cup_{i=1}^{n}\overline{D}_{i}))\subseteq\overline{D}_{0}.$
Proof.
Note that $F(\infty)=0\in\overline{D}_{0}$. Now let us pick $z_{0}\in K\setminus(\cup_{i=1}^{n}\overline{D}_{i})$ and estimate $|f_{i}(z_{0})h_{i}(z_{0})|$ for each $1\leq i\leq n$. Since $f_{i}$ is linear, we can write $f_{i}(z)=\alpha_{i}z+\beta_{i}$ for any $z\in\mathbb{C}_{K}$ for some $\alpha_{i}\in K\setminus\{0\}$ and $\beta_{i}\in K$. It follows that
$f_{i}(z_{0})h_{i}(z_{0})=\frac{\alpha_{i}z_{0}+\beta_{i}}{1-\left(\frac{z_{0}-a_{i}}{b_{i}}\right)^{M}}.$
From the choice of $b_{i}$, we have that
$\left|\frac{z_{0}-a_{i}}{b_{i}}\right|\geq\frac{1}{|\sqrt{\mathfrak{u}}|}>1.$
Thus
$|f_{i}(z_{0})h_{i}(z_{0})|=|\alpha_{i}z_{0}+\beta_{i}|\left|\frac{b_{i}}{z_{0}-a_{i}}\right|^{M}.$
If $|\alpha_{i}z_{0}|\leq|\beta_{i}|$, then
$|f_{i}(z_{0})h_{i}(z_{0})|\leq|\beta_{i}||\sqrt{\mathfrak{u}}|^{M}.$
Thus there exists $M_{i}^{(0)}\geq 1$ such that $|f_{i}(z_{0})h_{i}(z_{0})|\leq|\mathfrak{u}|$ for any $M\geq M_{i}^{(0)}$.
We now consider the case where $|\alpha_{i}z_{0}|>|\beta_{i}|$. Then
$|f_{i}(z_{0})h_{i}(z_{0})|\leq|\alpha_{i}z_{0}|\left|\frac{b_{i}}{z_{0}-a_{i}}\right|^{M}=\left|\frac{\alpha_{i}z_{0}}{z_{0}-a_{i}}\right||b_{i}|\left|\frac{b_{i}}{z_{0}-a_{i}}\right|^{M-1}\leq\left|\frac{\alpha_{i}z_{0}}{z_{0}-a_{i}}\right||b_{i}||\sqrt{\mathfrak{u}}|^{M-1}.$
If $|z_{0}|>|a_{i}|$, then $|z_{0}-a_{i}|\leq|z_{0}|$ and hence
$|f_{i}(z_{0})h_{i}(z_{0})|\leq|\alpha_{i}b_{i}||\sqrt{\mathfrak{u}}|^{M-1}.$
If $|z_{0}|<|a_{i}|$, then $|z_{0}-a_{i}|=|a_{i}|$ and hence $|(\alpha_{i}z_{0})/(z_{0}-a_{i})|<|\alpha_{i}|$; so we also obtain that
$|f_{i}(z_{0})h_{i}(z_{0})|\leq|\alpha_{i}b_{i}||\sqrt{\mathfrak{u}}|^{M-1}.$
If $|z_{0}|=|a_{i}|$, since $z_{0}\not\in\overline{D}_{i}$, we have that $|z_{0}-a_{i}|\geq\mathrm{diam}(\overline{D}_{i})$; so we obtain that
$|f_{i}(z_{0})h_{i}(z_{0})|\leq\frac{\left|\alpha_{i}a_{i}b_{i}\right|}{\mathrm{diam}(\overline{D}_{i})}|\sqrt{\mathfrak{u}}|^{M-1}.$
Thus in this case, there exist $M_{i}^{(1)}\geq 1$ such that $|f_{i}(z_{0})h_{i}(z_{0})|\leq|\mathfrak{u}|$ for any $M\geq M_{i}^{(1)}$.
Set $M_{i}=\max\{M_{i}^{(0)},M_{i}^{(1)}\}$. Then $|f_{i}(z_{0})h_{i}(z_{0})|\leq|\mathfrak{u}|$ for any $M\geq M_{i}$. Now we take
$M\geq\max\{M_{i}:1\leq i\leq n\}$
to be an even integer and conclude that, for any $z_{0}\in\mathbb{P}^{1}(K)\setminus(\cup_{i=1}^{n}\overline{D}_{i})$,
$|F(z_{0})|\leq\max\{|f_{i}(z_{0})h_{i}(z_{0})|:1\leq i\leq n\}\leq|\mathfrak{u}|.$
Thus the conclusion holds. ∎
We now pick $M$ as in Lemma 4.11. Since $F(\overline{D}_{0})=f_{0}(\overline{D}_{0})=\overline{D}_{0}$, we have that $\overline{D}_{0}$ is contained in the Fatou set of $F$. Then $\mathbb{P}^{1}(K)\setminus(\cup_{i=1}^{n}\overline{D}_{i})$ is also contained in the Fatou set of $F$ by Lemma 4.11. Hence
$\mathcal{J}(F,K)\subseteq\cup_{i=1}^{n}\overline{D}_{i}.$
Lemma 4.12.
The Julia set $\mathcal{J}(F,K)$ intersects $\overline{D}_{i}$ for each $1\leq i\leq n$.
Proof.
Pick any $1\leq i_{0}\leq n$. By Lemma 4.8 (2), we have that $F(\overline{D}_{i_{0}})\setminus(\cup_{i=1}^{n}\overline{D}_{i})\not=\emptyset$. Then by Lemma 4.11, we conclude that
$F^{\circ 2}(\overline{D}_{i_{0}})\cap\overline{D}_{0}\not=\emptyset.$
Again by Lemma 4.8 (2), we have that
$F^{\circ 2}(\overline{D}_{i_{0}})\cap(\cup_{i=1}^{n}\overline{D}_{i})\not=\emptyset.$
Thus there exist $x_{i_{0}},y_{i_{0}}\in\overline{D}_{i_{0}}$ such that $|F^{\circ 2}(x_{i_{0}})-F^{\circ 2}(y_{i_{0}})|\geq 1$.
Now given $1\leq j\leq n$, for any $m\geq 1$, consider a closed disk $\overline{D}_{j_{m}}\subseteq F^{\circ m}(\overline{D}_{j})$ with $1\leq j_{m}\leq n$. Let $\overline{D}_{j,m}\subseteq\overline{D}_{j}$ be a closed disk where $F^{\circ m}$ is bijective to $\overline{D}_{j_{m}}$. The existence of such a $\overline{D}_{j,m}$ follows from the fact that $F$ is linear in each $\overline{D}_{i}$ for $1\leq i\leq n$. Then passing to a subsequence in $m$, we obtain that as $m\to\infty$,
$\mathrm{diam}(\overline{D}_{j,m})\to 0.$
Consider $F^{\circ m}$-preimages $x_{j,m},y_{j,m}$ in $\overline{D}_{j,m}$ for the points $x_{j_{m}},y_{j_{m}}$ in $\overline{D}_{j_{m}}$, respectively. We conclude that, as $m\to\infty$,
$|x_{j,m}-y_{j,m}|\to 0,$
but
$|F^{\circ{(m+2)}}(x_{j,m})-F^{\circ{(m+2)}}(y_{j,m})|=|F^{\circ 2}(x_{j_{m}})-F^{\circ 2}(y_{j_{m}})|\geq 1.$
Thus $\{F^{\circ n}\}_{n\geq 1}$ is not equicontinuous in $\overline{D}_{j}$, equivalently, $\mathcal{J}(F,K)\cap\overline{D}_{j}\not=\emptyset$. Hence the conclusion holds. ∎
Proof of Proposition 4.10.
By Lemmas 4.8 and 4.12, it suffices to check the hyperbolicity of $F$. By Lemma 4.9 (2) and the linearity of $f_{i},1\leq i\leq n$, we conclude that in each $\overline{D}_{i}$ the map $F$ has degree $1$ and hence has no critical points. So $F$ has no critical points in $\mathcal{J}(F,K)$. Thus, $F$ is hyperbolic. ∎
4.2.5. Proof of Theorem 1.2 and examples
Proof of Theorem 1.2.
Let $\lambda\geq 1$ be a weak Perron number. If $\lambda=1$, we consider a rational map $\phi\in K(z)$ of degree at least $2$ having a good reduction. Then $J(\phi,K)=\emptyset$ and hence $\phi$ is a hyperbolic map with $\mathfrak{h}(\phi,K)=0=\log 1$. If $\lambda>1$, by Proposition 4.5, we consider an admissible $(0,1)$-matrix $A$ with leading eigenvalue $\lambda$. Then the map $\phi:=F$ in Proposition 4.10 is a hyperbolic rational map in $K(z)$. Moreover, by Proposition 3.9, $(\mathcal{J}(\phi,K),\phi)$ is topologically conjugate to $(\Sigma_{A},\sigma)$, which implies that $\mathfrak{h}(\phi,K)=\log\lambda$.
Conversely, let $\phi\in K(z)$ be a subhyperbolic rational map in $K(z)$ with compact $\mathcal{J}(\phi,K)$ such that $\mathfrak{h}(\phi,K)\not=0$. Then there exist a matrix $B$ with non-negative entries and finitely many numbers $\ell_{1},\dots,\ell_{m_{0}}$ such that
$\zeta_{\phi}(K,t)=\frac{\prod_{m=1}^{m_{0}}(1-t^{\ell_{m}})}{\det(I-tB)},$
see §3.5. Noting that the numerator and denominator of $\zeta_{\phi}(K,t)$ have no common zeros differing from roots of unity, we conclude that the reciprocal of the convergence radius of $\zeta_{\phi}(K,t)$ is a weak Perron number $\lambda_{B}>1$; in fact, $\lambda_{B}$ is the leading eigenvalue of $B$. Hence $\mathfrak{h}(\phi,K)=\log\lambda_{B}$. ∎
Example 4.13 (A hyperbolic rational map $\phi\in\mathbb{Q}_{2}(z)$ with $\exp(\mathfrak{h}(\phi,\mathbb{Q}_{2})=(1+\sqrt{5})/2$).
Let $A=\begin{bmatrix}1&1\\ 1&0\end{bmatrix}.$ Then $A$ is an admissible (0,1)-matrix with leading eigenvalue $(1+\sqrt{5})/2$. Let us consider the disks $\overline{D}(1,|8|)$ and $\overline{D}(3,|4|)$ in $\mathbb{Q}_{2}$, and define the maps
  • $f_{1}:\overline{D}(1,|8|)\to\overline{D}(1,|2|)$, sending $z\in\overline{D}(1,|8|)$ to $(z+3)/4\in\overline{D}(1,|2|)$, and
  • $f_{2}:\overline{D}(3,|4|)\to\overline{D}(1,|4|)$, sending $z\in\overline{D}(3,|4|)$ to $z-2\in\overline{D}(1,|4|)$.
Moreover, let $f_{0}:\overline{D}(0,|2|)\to\overline{D}(0,|2|)$ be the identity map. See Figure 2.
Figure 2. The linear maps $f_{1},f_{2}$ and $f_{0}$ in Example 4.13.
Computing a large even integer $M$ as in the proof of Lemma 4.11, we can take $M=14$. Then we obtain a desired hyperbolic rational map $\phi\in\mathbb{Q}_{2}(z)$ defined by
$\phi(z)$ $=\frac{f_{1}(z)}{1-\left(\frac{z-1}{4\sqrt{2}}\right)^{14}}+\frac{f_{2}(z)}{1-\left(\frac{z-3}{2\sqrt{2}}\right)^{14}}+\frac{f_{0}(z)}{1-\left(\frac{z}{\sqrt{2}}\right)^{14}}$
$=\frac{2^{33}(z+3)}{2^{35}-(z-1)^{14}}+\frac{2^{21}(z-2)}{2^{21}-(z-3)^{14}}+\frac{2^{7}z}{2^{7}-z^{14}}.$
Example 4.14 (A hyperbolic rational map $\phi\in\mathbb{Q}_{2}(z)$ with $\mathfrak{h}(\phi,\mathbb{Q}_{2})=0$ and $\mathcal{J}(\phi,\mathbb{Q}_{2})\not=\emptyset$).
Let $A=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}.$ Then $A$ is an admissible (0,1)-matrix with leading eigenvalue $1$. Let us consider the disks $\overline{D}(1,|8|)$ and $\overline{D}(3,|8|)$ in $\mathbb{Q}_{2}$, and define the maps
  • $f_{1}:\overline{D}(1,|8|)\to\overline{D}(3,|4|)$, sending $z\in\overline{D}(1,|8|)$ to $(z+5)/2\in\overline{D}(3,|4|)$, and
  • $f_{2}:\overline{D}(3,|8|)\to\overline{D}(1,|4|)$, sending $z\in\overline{D}(3,|8|)$ to $(z-1)/2\in\overline{D}(1,|4|)$.
Moreover, let $f_{0}:\overline{D}(0,|2|)\to\overline{D}(0,|2|)$ be the identity map. See Figure 3.
Figure 3. The linear maps $f_{1},f_{2}$ and $f_{0}$ in Example 4.14.
Computing a large even integer $M$ as in the proof of Lemma 4.11, we can take $M=6$. Then we obtain a hyperbolic rational map $\phi\in\mathbb{Q}_{2}(z)$ defined by
$\phi(z)$ $=\frac{f_{1}(z)}{1-\left(\frac{z-1}{4\sqrt{2}}\right)^{6}}+\frac{f_{2}(z)}{1-\left(\frac{z-3}{4\sqrt{2}}\right)^{6}}+\frac{f_{0}(z)}{1-\left(\frac{z}{\sqrt{2}}\right)^{6}}$
$=\frac{2^{14}(z+5)}{2^{15}-(z-1)^{6}}+\frac{2^{14}(z-1)}{2^{15}-(z-3)^{6}}+\frac{2^{3}z}{2^{3}-z^{6}}.$
In $\overline{D}(1,|8|)$, since $\phi(\overline{D}(3,|4|)\setminus\overline{D}(3,|8|))\subseteq\overline{D}(0,|2|)$, we have that
$\overline{D}(1,|8|)\subsetneq\phi^{\circ 2}(\overline{D}(1,|8|))\subseteq\overline{D}(1,|4|)\cup\overline{D}(0,|2|).$
Note that $\phi^{\circ 2}$ is a scaling map in $\overline{D}:=\phi^{-2}(\overline{D}(1,|4|))\cap\overline{D}(1,|8|)$. Thus $\overline{D}$ contains a fixed point, say $x_{0}$, of $\phi^{\circ 2}$. Since
$\phi^{\circ 2}(\overline{D}(1,|4|)\setminus\overline{D}(1,|8|))\subseteq\overline{D}(0,|2|),$
we conclude that any point in $\overline{D}(1,|8|)\setminus\{x_{0}\}$ is eventually mapped into $\overline{D}(0,|2|)$ under $\phi$ and hence is contained in $\mathcal{F}(\phi,\mathbb{Q}_{2})$. Thus $\mathcal{J}(\phi,\mathbb{Q}_{2})$ coincides with the two-cycle $\{x_{0},\phi(x_{0})\}$.
5. Example
In this section, we compute a concrete example for the Artin-Mazur zeta function and the entropy.
Given any prime number $p\geq 2$, let $K=\mathbb{Q}_{p}^{un}$ be the maximal unramified extension of $\mathbb{Q}_{p}$. Consider the polynomial $\phi\in K[z]$, acting as a self-map on $\mathbb{P}^{1}(K)$, defined by
$\phi(z)=\frac{z(z-1)^{2}}{8p}.$
We have the following observation:
  • the critical points of $\phi$ in $K$ are $1$ and $1/3$;
  • the point 00 is a repelling fixed point and hence $0\in\mathcal{J}(\phi,K)$;
  • $\phi(1)=0$ and hence $1\in\mathcal{J}(\phi,K)$;
  • $\phi^{\circ n}(1/3)\to\infty$, as $n\to\infty$, and hence $1/3\in\mathcal{F}(\phi,K)$; and
  • $\phi^{\circ n}(z)\to\infty$, as $n\to\infty$, for any $z\not\in D(0,1)\cup D(1,1)$.
In particular, $\phi$ is subhyperbolic. Moreover, $\mathcal{J}(\phi,K)$ is compact by [17, Theorem 4.8].
We now compute $\zeta_{\phi}(K,t)$ and $\mathfrak{h}(\phi,K)$ depending on $p\not=2$ or $p=2$. In the former case, the critical point $1$ is tame, and in the latter case, the critical point $1$ is wild.
Case 1: $p\not=2$. For each $n\geq 1$, consider the sets $A_{n}:=\overline{D}(0,|p^{n}|)\setminus D(0,|p^{n}|)$ and $\tilde{A}_{n}:=\overline{D}(1,|p^{n}|)\setminus D(1,|p^{n}|)$
In $A_{n}$, since $\phi(\overline{D}(0,|p|))=\overline{D}(0,1)=\overline{D}(1,1)$ and $\phi$ has degree $1$ on $\overline{D}(0,|p|)$, we conclude that
  • in $A_{1}$, there is unique closed disk $\overline{D}_{1}$ such that $\phi(\overline{D}_{1})=\overline{D}(1,|p|)$;
  • for each $n\geq 2$, in $A_{n}$, there is unique closed disk $\overline{D}_{n}$ such that $\phi(\overline{D}_{n})=\overline{D}_{n-1}$;
  • for each $n\geq 1$, $\phi^{\circ(n+1)}(A_{n}\setminus\overline{D}_{n})=K\setminus(D(0,1)\cup D(1,1))$ and hence $A_{n}\setminus D_{n}\subseteq\mathcal{F}(\phi,K)$.
It follows that
$D(0,1)\cap\mathcal{J}(\phi,K)\subseteq\bigcup_{n\geq 1}\overline{D}_{n}\cup\{0\}.$
In $\tilde{A}_{n}$, since $\phi(\overline{D}(1,|p^{n}|))=\overline{D}(0,|p^{2n-1}|)$ and $\phi$ has degree $2$ on $\overline{D}(1,|p^{2n}|)$, we conclude that
  • for each $n\geq 1$, in $\tilde{A}_{n}$, there are exactly two distinct closed disk $\overline{B}_{n}^{+}$ and $\overline{B}_{n}^{-}$ such that $\phi(\overline{B}_{n}^{\pm})=\overline{D}_{2n-1}$.
It follows that
$D(1,1)\cap\mathcal{J}(\phi,K)\subseteq\bigcup_{n\geq 1}(\overline{B}_{n}^{+}\cup\overline{B}_{n}^{-})\cup\{1\}.$
Thus
$\mathcal{J}(\phi,K)\subseteq\bigcup_{n\geq 1}(\overline{D}_{n}\cup\overline{B}_{n}^{+}\cup\overline{B}_{n}^{-})\cup\{1\}\cup\{0\}.$
Moreover, we obtain countable symbols $\overline{D}_{n}$, $\overline{B}_{n}^{+}$ and $\overline{B}_{n}^{-}$ (see Figure 4).
Figure 4. The portrait of $\overline{D}_{n}$, $\overline{B}_{n}^{+}$ and $\overline{B}_{n}^{-}$ under $\phi$.
We now identify $\overline{D}_{2n-1},n\geq 1$ as a symbol $\alpha^{+}$, $\overline{D}_{2n},n\geq 1$ as a symbol $\alpha^{-}$, $\overline{B}_{n}^{+},n\geq 1$ as a symbol $\beta^{+}$, and $\overline{B}_{n}^{-},n\geq 1$ as a symbol $\beta^{-}$, exchanging the disks $\overline{B}_{n}^{-}$ and $\overline{B}_{n}^{+}$ for some $n$ if necessary. Then adding the critical point $\{1\}$ and the critical value $\{0\}$, we obtain a finite graph (see Figure 5) with transition matrix $A$ given by
[…]
Figure 5. A finite graph for $\phi$ when $p\not=2$.
Observe that the two-cycle of periodic points $(\alpha^{+},\alpha^{-},\alpha^{+},\alpha^{-},\dots)$ and $(\alpha^{-},\alpha^{+},\alpha^{-},\alpha^{+},\dots)$ in the symbolic space $(\Sigma_{A},\sigma)$ induced by the above symbols is the only periodic cycle not corresponding to a periodic cycle in $\mathcal{J}(\phi,K)$. Thus we conclude that
$\zeta_{\phi}(K,t):=\frac{1-t^{2}}{\det(I-tA)}=\frac{1-t^{2}}{(1-t)(1-3t^{2})}=\frac{1+t}{1-3t^{2}}.$
Hence $\mathfrak{h}(\phi,K)=\log\sqrt{3}$.
Case 2: $p=2$. For each $n\geq 1$, consider the sets $A_{n}:=D(0,|p^{4n-4}|)\setminus D(0,|p^{4n}|)$ and $\tilde{A}_{n}:=D(1,|p^{2n-2}|)\setminus D(1,|p^{2n}|)$.
In $A_{n}$, since $\phi(\overline{D}(0,|p^{4}|))=\overline{D}(0,1)=\overline{D}(1,1)$ and $\phi$ has degree $1$ on $\overline{D}(0,|p^{4}|)$, we conclude that
  • in $A_{1}$, there are unique closed disks $\overline{D}_{1}\subseteq\overline{U}_{1}$ such that $\phi(\overline{D}_{1})=\overline{D}(1,|p^{3}|)$ and $\phi(\overline{U}_{1})=\overline{D}(1,|p^{2}|)$;
  • for each $n\geq 2$, in $A_{n}$, there are unique closed disks $\overline{D}_{n}\subset\overline{U}_{n}$ such that $\phi(\overline{D}_{n})=\overline{D}_{n-1}$ and $\phi(\overline{U}_{n})=\overline{U}_{n-1}$;
  • for each $n\geq 1$, $\phi^{\circ(n+1)}(A_{n}\setminus\overline{U}_{n})=K\setminus(D(0,1)\cup D(1,1))$ and hence $A_{n}\setminus U_{n}\subseteq\mathcal{F}(\phi,K)$.
It follows that
$D(0,1)\cap\mathcal{J}(\phi,K)\subseteq\bigcup_{n\geq 1}\overline{U}_{n}\cup\{0\}.$
In $\tilde{A}_{n}$, since $\phi(\overline{D}(1,|p^{2n}|))=\overline{D}(0,|p^{4n-4}|)$ and $\phi$ has degree $2$ on $\overline{D}(1,|p^{2n}|)$, we conclude that
  • for each $n\geq 2$, in $\tilde{A}_{n}$, there are exactly two distinct closed disks $\overline{B}_{n}^{+}$ and $\overline{B}_{n}^{-}$ such that $\phi(\overline{B}_{n}^{\pm})=\overline{D}_{n-1}$;
  • in $\tilde{A}_{1}$, there are exactly two distinct closed disks $\overline{B}_{1}^{+}$ and $\overline{B}_{1}^{-}$ such that $\phi(\overline{B}_{1}^{\pm})=\overline{D}(1,|p^{3}|)$;
  • in $\tilde{A}_{1}$, there are exactly two distinct closed disks $\overline{V}^{+}$ and $\overline{V}^{-}$ such that $\overline{B}_{1}^{\pm}\subseteq\phi(\overline{V}_{1}^{\pm})$ and $\mathrm{diam}(\phi(\overline{V}^{\pm}))=p\cdot\mathrm{diam}(\overline{B}_{1}^{\pm})$;
It follows that
$\overline{V}^{\pm}\subsetneq\phi(\overline{V}^{+})=\phi(\overline{V}^{-}).$
Moreover,
$\tilde{A}_{1}\cap\mathcal{J}(\phi,K)\subseteq\overline{B}_{1}^{+}\cup\overline{B}_{1}^{-}\cup\overline{V}^{+}\cup\overline{V}^{-}.$
In $\overline{U}_{n}$, since $\phi(\overline{U}_{1})=\overline{D}(1,|p^{2}|)$, we conclude that
  • in $\overline{U}_{1}$, there is a unique closed disk $\overline{W}_{1}\not=\overline{D}_{1}$ such that $\phi(\overline{W}_{1})=\phi(\overline{V}_{1}^{\pm})$;
  • for each $n\geq 2$, in $\overline{U}_{n}$, there is a unique closed disk $\overline{W}_{n}\not=\overline{D}_{n}$ such that $\phi(\overline{W}_{n})=\overline{W}_{n-1}$;
Then in $\tilde{A}_{n}$, we conclude that
  • for each $n\geq 2$, in $\tilde{A}_{n}$, there are exactly two distinct closed disks $\overline{X}_{n}^{+}$ and $\overline{X}_{n}^{-}$, distinct from $\overline{B}_{n}^{\pm}$ such that $\phi(\overline{X}_{n}^{\pm})=\overline{W}_{n-1}$.
It is easy to check that
$\mathcal{J}(\phi,K)\subseteq\bigcup_{n\geq 1}(\overline{D}_{n}\cup\overline{W}_{n}\cup\overline{B}_{n}^{+}\cup\overline{B}_{n}^{-}\cup\overline{X}_{n}^{+}\cup\overline{X}_{n}^{-})\cup\overline{V}^{+}\cup\overline{V}^{-}\cup\{1\}\cup\{0\}.$
Moreover, we obtain countable symbols $\overline{D}_{n},\overline{W}_{n}$, $\overline{B}_{n}^{\pm}$,$\overline{X}_{n}^{\pm}$ and $\overline{V}^{\pm}$ (see Figure 6).
Figure 6. The portrait of $\overline{D}_{n},\overline{W}_{n}$, $\overline{B}_{n}^{\pm}$,$\overline{X}_{n}^{\pm}$ and $\overline{V}^{\pm}$ under $\phi$.
We now identify $\overline{D}_{n},n\geq 1$ as a symbol $\alpha$, $\overline{W}_{n},n\geq 1$ as a symbol $\omega$, $\overline{B}_{n}^{+},n\geq 2$ as a symbol $\beta^{+}$, $\overline{B}_{n}^{-},n\geq 2$ as a symbol $\beta^{-}$, $\overline{B}_{1}^{+}$ as a symbol $\beta_{1}^{+}$, $\overline{B}_{1}^{-}$ as a symbol $\beta_{1}^{-}$, $\overline{X}_{n}^{+},n\geq 2$ as a symbol $\theta^{+}$, $\overline{X}_{n}^{-},n\geq 2$ as a symbol $\theta^{-}$, $\overline{V}^{+}$ as a symbol $v^{+}$, $\overline{V}^{-}$ as a symbol $v^{-}$, exchanging the disks $\overline{B}_{n}^{-}$ and $\overline{B}_{n}^{+}$ for some $n$ and exchanging the disks $\overline{X}_{m}^{-}$ and $\overline{X}_{m}^{+}$ for some $m$ if necessary. Then adding the critical point $\{1\}$ and the critical value $\{0\}$, we obtain a finite graph (see Figure 7) with transition matrix $A$ given by
[…]
Figure 7. A finite graph for $\phi$ when $p=2$.
Observe that the two fixed points $(\alpha,\alpha,\alpha,\dots)$ and $(\omega,\omega,\omega,\dots)$ in the symbolic space $(\Sigma_{A},\sigma)$ induced by the above symbols are the only periodic cycles not corresponding to periodic cycles in $\mathcal{J}(\phi,K)$. Thus we conclude that
$\zeta_{\phi}(K,t):=\frac{(1-t)^{2}}{\det(I-tA)}=\frac{(1-t)^{2}}{(1-t)^{2}(1-3t)}=\frac{1}{1-3t}.$
Hence $\mathfrak{h}(\phi,K)=\log 3$.